### v1

# Establishing validated standards for Home and Work location Detection

Abstract
Smartphone location data have transformed urban mobility research, providing unprecedented insights into how people navigate and interact in cities. However, leveraging location data at scale presents methodological challenges. Accurately identifying individuals’ home and work locations is critical for a range of applications, including commuting analysis, unemployment estimation, and urban accessibility studies. Despite their widespread use, home–work detection methods lack a standardized framework that accounts for differing data quality and that is validated against ground-truth observations. This limits the comparability and reproducibility of results across studies and datasets. In this paper, we present HoWDe, a robust algorithm for identifying home and work locations from mobility data, explicitly designed to handle missing data and varying data quality across individuals. Using two unique ground-truth datasets comprising over $5,100$ individuals from more than $80$ countries, HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across countries and demographic groups. We examine how parameter choices shape the trade-off between accuracy and user retention, and demonstrate how these methodological decisions influence downstream applications such as employment estimation and commuting pattern analysis. By supporting in-house pre-processing through a transparent and validated pipeline, HoWDe also facilitates the sharing of privacy-preserving mobility data. Together, our tools and findings establish methodological standards that support more robust, scalable, and reproducible mobility research at both individual and urban scales.
Keywords: GPS data, Home and work detection, Human Mobility, Stop-location sequences, Mobility algorithms, Missing data, Urban analytics
1 Introduction
Location data collected from smartphones has become indispensable for human mobility analyses [1, 2]. These datasets allow studying large samples of individuals globally, providing insights on fundamental aspects of human movements [3], epidemic spreading [4, 5, 6, 7], migration patterns [8, 5], transportation modeling [9], urban planning [10, 11], and the impact of natural disasters and societal shocks on population movements [12, 13, 14, 15, 16]. In light of these advancements, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [17].
A critical component of location data analyses involves the detection of individuals’ visits to their residential [18, 19, 20], and work location [17]. The accurate identification of home and work locations is key for mobility research, as it enables the analysis of factors such as socio-economic background [21, 22, 15], commuting behaviors [23, 16], social contact networks [24, 14], and accessibility to amenities and services [25, 14]. More broadly, it contributes to advancements in computational demography and social science research [26, 17].
The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behavior, such as the tendency to spend nights at home and working hours at a workplace [27, 28, 29]. These methods face significant challenges. First, they are sensitive to missing data, which is common in smartphone records [17]. Data quality varies across individuals and over time, as location gaps arise due to factors like inactivity, disabled services, or poor signal, introducing biases in home and work inference [30, 31, 17]. Second, they rely on the choice of parameters, such as time span required to establish a home location or the minimum number of hours per day spent at work.
Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains unclear, and there is no consensus on how to choose one specific algorithm and its parameters’ values [21, 32, 33, 34, 35, 17], in part due to a lack of ground truth data. Validation efforts so far have been limited, focusing on small subsets of individuals, home detection alone, and datasets with low spatial resolution and short time periods [34, 35, 32, 21]. Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been understood in relation to algorithmic performance. The lack of thorough validation has hindered the refinement and harmonization of these methods into standardized best practices [36, 17]. In turn, the absence of such standards can impact the comparability and reproducibility of mobility studies [17].
In this work, we leverage two distinct ground truth datasets to address these key challenges in home/work location detection. Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 (see Methods for details).
First, we leverage these data to uncover behavioral patterns in home and work visits. This allows us to empirically assess whether the assumptions underlying home and work detection hold in practice. Additionally, it enables us to quantify the variability in home-work visit behavior across a diverse set of users.
Second, we formalize HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach. HoWDe combines the explicit detection of both home and work locations [14, 29], the identification of the workplace based on recurrent visits during working hours [29], the identification of the home as the most frequently visited location at night [35, 28, 34, 21], a sliding window to identify home and job changes over time [14], with tunable window length [14]. Additionally, HoWDe improves upon existing approaches by better accounting for data gaps and their variation across users and over time. We provide an open-source implementation of the algorithm [37].
Finally, we assess HoWDe’s performance against the ground truth data. We provide a detailed analysis of the role of its parameters, and we study how performance changes across age groups, gender, countries, and urbanization levels. We conclude by quantifying how errors in the estimation of home and work locations propagate to key downstream tasks such as estimating employment rates or commuting distances.
2 Results
2.1 Unveiling temporal patterns of home-work visits
We analyze the regularity of home and work visits by measuring how often and when individuals are recorded at their annotated locations. Figure 1a-b shows the probability of being at home or work by hour of the day. We find that home visits peak between midnight and 6 AM, while work visits are most frequent during typical business hours (between 9 AM and 4 PM).
Patterns of visits to the home locations are similar between weekdays and weekends (see Figure 1c-e). For the work location, we observe a drastic reduction in the fraction of individual visiting their work location during weekends (see Figure 1d-f).
Results are consistent across both datasets, with overall higher probabilities of being at home (and lower probabilities of being at work) found in D2 due to the COVID-19 pandemic containment policies in place during most of the period covered by the data (see Sec.SI-3).
Our analysis suggests that work location is best determined using weekday daytime hours, while home location can be identified using night hours, regardless of weekdays or weekends, similar to previous approaches [2, 28, 35].
However, while general trends exist, we also find significant variations across days and users, highlighting pronounced behavioral heterogeneities at the individual level. Figure 1c-f reports disaggregated visit patterns for single “user-day” pairs. Most users’ daily patterns can be grouped into six behavioral profiles based on home and three on workday visits (see Methods for details on how behavioral profiles are obtained).
Figure 1: Behavioral visit profiles to home and work locations a. Fraction of visits per hour of the day to the annotated home locations in dataset D1, for weekdays (black) and weekends (grey) across individuals. The shaded orange area highlights the period with the highest probability of being at home. b. Fraction of visits per hour of the day to the annotated work locations in dataset 1, for weekdays and weekends. The shaded blue area highlights the period with the highest probability of being at work during weekdays. c. User-day profiles by hours at home (orange), other locations (grey), and periods without location data (light-grey) during weekdays. d. Hours at work (blue), other location (grey), or without data (light-grey), per hour of the day across days with data for weekdays. e-f. Show user-day profiles during weekends for hours at home (orange) and work (blue) respectively. The user-day profiles are sorted by cluster, hour of the first home visit, total time spent at home, and hours without data. Clusters are separated by black vertical lines, and the fraction of days per cluster is noted on the x-axis.
With respect to home visits, most user days fall into one of three typical profiles: i) days when users mostly stay at home, ii) when they are rarely/never at home, and iii) when they are away from home during typical business hours (see Figure 1c,e). Similarly, work-behavior profiles are described by three main daily patterns: i) days when users spend business hours at work, ii) when they are rarely/never at work, and iii) when they spend night hours at work (Figure 1d,f). We find that the distribution of individuals across these profiles is generally independent of their socio-demographic characteristics, such as age, gender, and urbanization (see Sec.SI-5).
As the behavioral profiles we identified are based on single user-day activity data, an important question is whether individuals switch between these profiles or, rather, a single profile consistently represents them. We find that users visit a median of $5$ out of $6$ home clusters and $2$ out of $3$ work clusters (for both D1 and D2). We measure the individual inter-profile variability through the normalized entropy, a measure ranging from 00 (when all days of a user fall into a single profile) to $1$ (when all profiles are equally important) [38] (see Methods). On average, users exhibit a normalized entropy of $0.66\pm 0.01$ for home-profile changes between weekdays and $0.48\pm 0.01$ for weekends ($0.62\pm 0.01$ and $0.42\pm 0.01$ for D2 respectively), and $0.51\pm 0.01$ for work-profile changes on weekdays only ($0.40\pm 0.02$ for D2). These values are comparable to the ones obtained when randomly allocating profiles among user days, confirming that users do not have a single dominant behavioral profile (see Table si.1).
All in all, our findings suggest that focusing on the same profiles across users is unlikely to favour one specific socio-demographic group, thus supporting the choice of focusing on weekday daytime hours (9 am to 4 pm) for detecting work locations and nighttime hours (12 am to 6 am) for detecting home locations, regardless of weekdays and weekends. These time windows, combined with minimum data coverage requirements, form the foundation of the heuristics implemented by the HoWDe algorithm.
2.2 HoWDe: a Home and Work Detection algorithm
In this section, we introduce HoWDe, a dynamic home and work detection algorithm that can capture residential moves and job changes over time through a sliding window approach. A key feature of HoWDe is the identification of home and work locations based on the proportion of time spent at a given location, rather than the absolute amount of time as approached in previous literature [29, 28, 14, 15, 33, 34]. This enables HoWDe to accommodate users with varying amounts of missing data over time (see Figure 1c-f), through tunable parameters that set data-quality filters. We share an open-source implementation of the algorithm in PySpark [37].
2.2.1 HoWDe workflow
HoWDe takes as input a user’s pre-processed sequence of stops, where each stop is fully defined by a user-location unique ID and a start and end timestamp.
First, we perform a temporal aggregation of the stop sequences into hourly-bin sequences. In each hourly bin, we retain the location where the individual has spent most of the time. This aggregation limits sequences to a maximum of 24 locations per day, boosting computational efficiency and making HoWDe faster than previous approaches based on sliding windows, with only limited information loss [14, 15].
Second, we exclude days with insufficient hourly bins with data (this step is controlled by the tunable parameter $dhn$, see Table 2). This is done separately for work detection (where we consider business hours) and home detection (where we consider night hours, see Figure 2b). Hence, it is possible that one day is retained only for the home detection or for the work detection.
Third, for each day $t$, we calculate the fraction of hourly-bins spent at each location during nighttime (24 pm to 6 am for home detection), and typical business hours (9 am to 4 pm on weekdays, for work detection). To compute such a fraction, we use as a denominator the number of hourly bins with records (Figure 2.b). Unlike previous approaches that rely on absolute time spent at each location, considering fractions can better accommodate users with varying amounts of missing data over time.
Fourth, for each day $t$, we consider a window of size $\Delta_{T}+1$, covering the range included in $t\pm\Delta_{T}/2$ days (see Figure 2.c). This window acts as a memory component for the algorithm to better detect behavior in the period before and after a specific day. In each window, we average the fraction of hours a location was visited (see Figure 2c).
Thus, for each window, we obtain a collection of potential home and work locations. Among these, we retain only those that meet minimum requirements with respect to the proportion of visits within and across days. These requirements are controlled by parameters $dn_{H,W}$, $hf_{H,W}$, $df_{H,W}$ (see Table 2). If no location meets the requirements, no home (or work) is assigned for that specific window.
Finally, we sort the home and work collections in two different ways. For home, we sort based on the typical fraction of hourly bins the location is visited within a day. For work, we sort based on the typical fraction of days the location is visited within the window. If no work location can be selected based on the fraction of days, the sorting is based on the fraction of hourly bins. The top location from these two lists is then assigned as home (or work). By prioritizing the fraction of days over the fraction of hours for work detection, we are emphasizing work as a location defined by regular visits throughout the workweek (see Sec.SI-6).
In summary, the algorithm allows the control of two aspects. It enables the user to specify the criteria for determining when there is enough data to attempt home and work detection. Also, it enables to control of when a location can be considered as a candidate home or work location, all while accounting for the fact that users can have different levels of missing data.
Figure 2: Schematics of the HoWDe algorithm workflow for home/work location detection. a. Network of visited locations. The diameter represents the number of visits. In orange, we report the stops visited during night hours, and in blue are the ones visited during work hours. b. Sequences of locations visited hourly. The pie charts display the allocation of visits across locations during the night hours (orange) and work hours (blue) on each day. Note that on Saturday, 6/01, the night hours data was discarded since there were not enough hours with available data (exemplifying the filter tuned by the parameter $dhn$). In this minimal example, no location appears during both night and work hours; however, this may happen in real data. c. For each day, $t$ we consider a sliding window of $3$ consecutive days centered on $t$. For work detection, the sliding window excludes weekends. d. Aggregation step for the window centered on Friday, 5th. For each location, we compute the average fraction of time it was visited during night hours (orange) and during work hours (blue). Additionally, we compute the fraction of days the location is visited at least once during work hours. Locations are then sorted in descending order based on these fractions, and the top ones are selected as the estimated Home (L1) and Work (L2) locations on Friday.
2.3 Validation against ground truth data
To assess how well the procedure described above can capture individuals’ home and work locations, we measure i) the accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Methods for more details). We run the validation using the ground-truth datasets D1 and D2, described in Methods. In dataset D2, annotators reviewed complete stop sequences; therefore, the temporal dimension (changes of home and work locations) can not be considered. For this reason, for D2, we consider a single window covering the entire data period (see Sec.SI-7 for parameter details).
2.3.1 Algorithm performance across parameter choices
We find that sliding window length $\Delta_{T}$ and minimum home-time fraction $hf_{H}$ have the greatest impact on the accuracy of HoWDe’s home detection, while other parameters have a smaller effect (see Sec.SI-8).
Both the accuracy of the home location detection and the fraction of non-detected homes increase as the minimum home-time fraction threshold $hf_{H}$ increases (Figure 3a,b). In dataset D1, the fraction of non-detected homes increases with the window size, but accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks). This suggests that a window of $4$ weeks offers a good balance between high accuracy and user retention for home detection.
Figure 3: Home and Work validation across configurations a. Accuracy of the home location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night. For dataset D1, we explore how the results change for each sliding window increase (x-axis), while for dataset D2, we focus on a single sliding window (comprising the entire data period). b. Fraction of not detected of the home location detection for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend. c. Accuracy of the work location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours, and the minimum fraction of days an individual should visit the potential work across weekdays. For dataset D1, we explore how the results change for each sliding window increase, while for dataset D2, we keep the same sliding window size (comprising the entire period). d. Fraction of not detected of the work location detection for dataset D1 (first column) and dataset D2 (second column). The legend is shared with panel c.
Table 1 reports the results by focusing on two configuration setups: the minimum configuration, which maximizes accuracy while minimizing the number of non-detected users, and the maximum configuration, which prioritizes maximum accuracy at the cost of higher user non-detection rates. We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [28, 2] and TimeGeo [29] (see Methods for more details on these algorithms).
The maximum configuration is reached for $\Delta_{T}=28$ days window and $90\%$ minimum home-time fraction (see Sec.SI-7 for parameter details). With this choice of parameters, the algorithm achieves an accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 1). The minimum configuration is obtained for parameters $\Delta_{T}=28$ days window and $50\%$ home frequency (see Sec.SI-7 for parameter details). Here, the algorithm achieves $92.94\%\pm 0.49\%$ accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
All in all, we find that HoWDe achieves higher accuracy than previous heuristic methods [29, 2], partly because it excludes users with low-quality data (see Table 1).
Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T}=42$days ($t\pm 3$ weeks), after which accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
In the work location detection, two key parameters are the minimum fraction of hourly bins ${hf_{W}}$ and days ${df_{W}}$ required for a location to be selected as work. Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance accuracy in D1 but increase non-detection rates. For a larger value of ${hf_{W}}=60\%$, increasing ${df_{W}}$ enables significant accuracy gains. In contrast, for ${hf_{W}}=40\%$, changes in ${df_{W}}$ have a weaker effect. We find similar results for D2, when changing the minimum fraction of days ${hf_{W}}$ and fraction of hours ${hf_{W}}$.
HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Sec.SI-7 for parameter details), outperforming baseline methods (see Table 1). Similarly, for D2, HoWDe achieves high accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 1). This difference in performances may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
Accuracy HoWDe min HoWDe max Atlas TimeGeo
Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $81.96\pm 0.55$ $79.28\pm 0.78$
D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.77\pm 1.21$ $70.38\pm 1.12$
Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $60.32\pm 0.80$ $39.51\pm 0.92$
D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $69.71\pm 2.17$ $57.04\pm 2.40$
Fraction not detected HoWDe min HoWDe max Atlas TimeGeo
Home D1 $4.29\pm 0.38$ $38.75\pm 0.93$ $4.60\pm 0.31$ $\mathbf{0.09\pm 0.06}$
D2 $10.44\pm 1.02$ $39.94\pm 1.43$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $3.45\pm 0.31$ $\mathbf{1.86\pm 0.23}$
D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.67\pm 0.41$
Table 1: Evaluation of home and work detection across methods and datasets.
2.4 Robustness and biases across configurations
HoWDe’s accuracy comes at the cost of user retention. The exclusion of users may introduce biases toward specific demographic groups. To assess these biases, we measure accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanization levels, and country GDP. Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Sec.SI-7).
The largest accuracy gaps appear in suburban areas and middle-income countries: home detection drops by $3.39$ points in suburban areas (compared to urban areas) and work detection by $9.13$ in middle-income countries (compared to high-income countries). Despite accuracy deviations, home detection accuracy remains within $90\%$ $99\%$ and work detection within $77\%$ $89\%$ across all lower-accuracy subgroups (see Sec. SI-9), aligning with the observed range for D1 and D2.
Nevertheless, even minor detection errors can propagate, compromising the overall reliability of the results. To address this potential issue, we examine how tuning the detection algorithm parameters impacts results in two downstream applications: (i) employment rate estimates in Europe and (ii) commuting pattern characteristics.
2.4.1 Estimating regional employment rates
We estimate the regional employment rate by measuring the fraction of individuals with a detected work location, stable for a minimum of $30$ days. We then investigate how the employment estimates obtained from a “minimum” and “maximum” configurations differ and how much they deviate from official regional statistics [39]. This analysis covers $90$ regions across six European countries, each with a minimum of $1000$ individuals per region.
Figure 4.a shows that both configurations produce similar geographical patterns in regional employment. However, while both exhibit moderate to high Pearson correlation, the maximum configuration has a stronger correlation with reported employment rates but estimates fewer regions due to stricter data coverage thresholds. It also has a higher overall relative error (Figure 4.b, and Figure SI-12, except in Germany, where the trend reverses. This suggests that while the maximum configuration enhances individual accuracy, the minimum configuration may be more suitable for population estimates.
2.4.2 Home-Work Commuting distances
This case study examines whether commuting distance, measured as the great-circle distance between detected home and work locations, varies with population density and whether parameter configurations can influence results. We estimate urbanization levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [40].
Our analysis reveals consistent findings across configurations: commuting distances are longer in rural areas compared to urban areas, with an absolute difference of $3.61\pm 0.04$ km for the “minimum” and $3.85\pm 0.06$ km for the “maximum” configurations (see Figure 4.c). Additionally, commuting distances decrease with increasing population density (see Figure 4.d). However, the maximum configuration results in shorter commuting distances, with the most densely populated areas showing the largest differences between configurations. This highlights the need for robustness checks to ensure consistency across different configurations.
Figure 4: Applications for home and work detection a. Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest accuracy (right). b. Difference between the minimum (light teal) and maximum configurations (dark teal) by country, as the correlation between the reported and selected configuration (top), the relative error (bottom). c. Difference in commuting distance (km) between urban (dashed) and rural (dotted) by configuration. d. Difference in commuting distance (km) against the median and standard error of the population density bins ($\frac{people}{km^{2}}$).
3 Discussion
In this work, we contributed to the development of accurate and robust methods for home and work detection. Using two datasets (D1, D2) covering $\sim 5,100$ individuals from over $80$ countries, we characterized home/work visit patterns. We found that individuals are most likely at home between midnight and 6 AM ($\sim 75\%$ probability) and at work between 9 AM and 4 PM on workdays ($\sim 50\%$). However, visit patterns and missing data vary widely across users and over time, and even within individual users’ routines.
Building on these insights, we provided an open-source implementation of HoWDe, a home/work detection algorithm that integrates existing approaches while improving robustness against missing data. We validated HoWDe against ground truth datasets, achieving up to $\sim 97\%$ accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ accuracy in D1 and $\sim 82\%$ accuracy in D2 for work detection.
Our analysis highlights the trade-off between accuracy and retaining users with low-quality data. We identified two key parameters affecting performance: the window length considered and the minimum visit proportion required for classification. A window of at least $28$ days ensures a reliable home estimation, while work windows longer than $42$ days reduce user retention without significant accuracy gains.
We also revealed how the algorithm’s accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection). Although these accuracy gaps persist across configurations, two case studies show that prioritizing accuracy over user retention is preferable for individual-level tasks, whereas maximizing retention may be more valuable for population-level analyses.
The insights gained from this study, along with a formalized algorithm, can support research on commuting [15, 16], social contact networks [6], and access to amenities [14], with implications for areas such as computational social science, urban planning, and public health [41, 17]. Improving location data processing will enhance our understanding of both universal mobility laws and micro-level social dynamics [42].
Despite its strengths, HoWDe has key limitations. While it allows customization of detection windows and thresholds, more advanced methods may be required to fully capture mobility patterns that are collected by diverse GPS-equipped devices [43]. Its performance may also vary with lower temporal resolution or geographically limited datasets, particularly in underrepresented low-income regions with sparser mobile data. Future work should assess its applicability in rural and less densely populated areas [44].
Detecting home and work locations raises important ethical concerns, particularly regarding privacy [45, 46], as the geographical coordinates of these locations are highly identifying and can enable re-identification of individuals in mobility datasets. Currently, many data holders apply in-house heuristics to detect home and work, but these methods are often opaque, introducing methodological uncertainty. By providing a validated and open-source detection framework, our approach enables data holders to perform this sensitive step internally using a standardized and transparent pipeline. We hope this will facilitate the sharing of privacy-preserving datasets—for example, through the release of obfuscated or aggregated representations—while ensuring methodological consistency and scientific utility for downstream research.
In summary, by advancing methods for home and work detection, this study contributes to a more robust and scalable foundation for mobility research. By providing an open-source, validated solution, we hope to facilitate more reliable and ethical research, contributing to a deeper understanding of human mobility and its implications for society.
4 Material and Methods
4.1 Dataset 1: DTU stop data and self-reported annotations
The first dataset re-uses data collected during 2017-19 already used in previous publications [3], where stop locations were detected using the scalable stop-location detection algorithm InfoStop [47]. The stop location data captures the mobility patterns of 96,065 users between 25 and 64 years old for 81 countries across continents.
The dataset also contains self-recorded home and work labels for $2,309$ labelled home and $2,479$ work locations, collected across $4,599$ mobile phone and smart band users. Note that the labelled dataset is biased towards urban areas and higher- to middle-income countries. The spatial coordinates of locations are no longer available to preserve individuals’ privacy.
All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).
4.2 Dataset 2: Veraset Movements data and annotation
Anonymous raw GPS points are provided at a global level by the Veraset data company in the framework of the World Bank “Monitoring COVID-19 policy response through human mobility data” project. Veraset Movements data consists of raw GPS point information from devices all around the world. In contrast to the majority of GPS data providers, Verasets Movements are sourced from thousands of Software Development Kits (SDK) to reduce sample biases. It features coverage of about $5\%$ of the global population [48].
In our work, unique-device GPS points are processed and analyzed as a proxy of individuals’ mobility and patterns of visits. The data at our disposal spans over a period of $12$ months, starting from Jan 01, 2020, until Dec 31, 2020. In this work, multiple countries were analyzed to assess the impact of the pandemic and the policy responses across a vast spectrum of potentially different behaviors.
Dataset D2, in particular, consists of the GPS trajectories of $100$ distinct devices for each of the $5$ selected countries: Brazil, Colombia, Indonesia, Mexico, and the Philippines. For each country independently, we require $33\%$ of the users to be living in high-wealth neighbourhoods, $33\%$ of the users to be living in medium-wealth neighbourhoods, and the remaining part in low-wealth neighbourhoods. Wealth classification and users’ place of residence are drawn leveraging previous work on GPS mobility in middle-income countries [15]. In particular, each user is also required to have a minimum level of recorded activity, consisting of at least a stop record in $20\%$ in the days before the COVID-19 pandemic declaration and $20\%$ of days with at least a stop record over the entire data period.
This second dataset involves, for each user’s stop location sequences, the expert annotation through detailed inspection of location visits, satellite imagery, and local amenities, focusing on five middle-income countries (Brazil, Colombia, Indonesia, Mexico, and the Philippines) for 500 individuals, from which 287 have at least one work location annotated (see Sec. SI-1 detailed description of the annotation procedure). Given that annotators only labelled the complete longitudinal stop sequences, an explicit consideration of changes in home and work for each individual is not possible.
4.3 Home and Work locations detection algorithms: a primer
Smartphones use an in-built GPS sensor to estimate the device’s latitudes and longitude coordinates over time, at a frequency that varies depending on device usage. The device position is then collected by various location-based apps, such as navigation or social media apps. These data may then be used by location intelligence companies or researchers to study human mobility and its regularities, both at the individual and at the population level. Raw location data utilized for mobility analysis consists of timestamped sequences of (latitude, longitude) records associated with a unique user ID.
The first pre-processing step involved in the analysis is the so-called “stop-location detection” [47, 49]. Stop location detection consists of (i) identifying records that are captured when individuals stay within a small area for a sufficient amount of time (e.g. when they are at home, at the office, or in a restaurant); (ii) applying spatial clustering to all these “static records”, to identify a set of unique locations, i.e. those places where people spend time. After detecting the individuals’ stops-locations, individual trajectories can be summarized as sequences of stays. Each stay is associated with a start time (when the stay started), a duration (how long the stay lasted), and a location ID (where the stay happened).
A second critical processing step often involved in mobility studies is to identify key locations, such as individuals’ home and work locations, starting from the sequences of stays. To achieve this, researchers and practitioners often define heuristics based on the assumption that visits to home and work locations follow specific daily and weekly patterns. While a standard practice does not exist, in this work, we focus on two distinct and widely used state-of-the-art methods that exploit travel circadian rhythms and pattern regularities. These methods were designed and tested over the past decade with different approaches, in a collective scientific effort, separately leveraging diverse sets of mobile-phone-based data (e.g. Call Detail Records (CDRs), eXtended Detail Records, (XDRs), or GPS records) [21, 50, 51, 35] as well as activity-based travel surveys [36]. The main rationale is to rank locations based on the day of the week (weekdays vs. weekends) and the time of day (daytime vs. nighttime) when visits occur.
4.3.1 The Atlas approach
The method used in [2, 6, 52, 28] (hereinafter “Atlas” method) defines a home location as the most probable location visited on weekdays between 10 p.m. and 6 a.m., i.e. the stop location where a device was found spending most of the weekday nighttime. While the Atlas methods mostly focuses on home locations, its definition can be easily generalized to the case of work locations. Following a similar rationale, we define an Atlas work location as the most probable stop location visited during weekdays daytime.
4.3.2 The Time-Geo approach
In contrast with the definition introduced by the Atlas method, more complex heuristics are adopted in Jiang et al. [29] (hereinafter referred to as the “Time-geo” method), in particular for the case of identifying work locations. Time-geo ranks stop locations based on the frequency of visits during weekday nights (between 7 p.m. and 8 a.m.) and weekends and picks the most visited location as an individual’s home. The same logic is applied when detecting the work location, but now focusing on weekday work time boundaries (between 8 a.m. and 7 p.m.) and including additional requirements: i) the candidate’s work location is visited at least three times across the study period, and ii) the distance between the work candidate and an individual’s home location is greater than $500$ meters.
Note that both methods rely on the definition of stop location. Stop location definition might vary depending on the type of data analyzed and the specific characteristics of the data points, e.g. whether you use Call Detail Records (CDR), eXtended Detail Records (XDR), or GPS points. However, the principled choices these algorithms are founded on can naturally be extended to any of these different cases, making them easily generalizable.
In this work, stop locations are defined as spatiotemporally close sets of GPS points. All comparisons between home-work detection methods are always drawn from a set of stop locations constructed following identical procedures. For more information about stop location processing, we refer to the Supplementary Information (SI) Sec.SI-2 and the existing literature [47, 14].
4.4 Home and work visits behavioral clusters
Section 2.1 focused on the temporal attributes of users’ daily mobility patterns. To better highlight groups of users with similar behavioral features, we reported the results in Figure  1.c-f, sorting daily patterns based on the visit hour of the first home visit. behavioral groups were generated using the K-Modes clustering algorithm, a categorical adaptation of the traditional K-Means algorithm [53]. Here, we use daily sequences of 24 hourly bins with three possible stop categories. For example, in the case of home patterns, each sequence consists of either a home (category 1), a visit to “a different location”, e.g., work or recreational locations (category 2), or an hour with missing data (category 3). A similar approach is used for work patterns, where visits to work location are considered as “category 1”, while visits to home location are considered as “a different location” (category 2). The K-Modes algorithm requires as a parameter, the number of desired clusters to compute. In Figure 1.c,e we reported 6 clusters. In Figure 1.d,f we reported 3 and 4 clusters, respectively. Note that, while the number of clusters is the result of the selection of cluster number following the elbow method for each of the four plots, our purpose is to simply highlight the main different visit patterns. Following this idea, in Sec.SI-3, where we present results for the smaller D2 dataset, we manually selected three as the number of clusters used to report the hourly day-visit patterns.
4.5 Behavioral profile entropy
We quantify the degree to which users transition between behavioral profiles measuring the normalized Shannon entropy [38], treating each behavioral cluster as a distinct state, following:
$H_{\text{norm}}=-\frac{1}{\log{k}}\sum_{i=1}^{k}p_{i}\log{p_{i}},$ (1)
where $k$ is the total number of possible clusters, and $p_{i}$ is the probability of occurrence of each state, calculated as: $p_{i}=\frac{n_{i}}{N}$ where $n_{i}$ are the total days classified in cluster $i$ and $N$ the total number of days for a given user. Normalizing by the maximum entropy $\frac{1}{\log{k}}$, ensures that entropy remains comparable across cases (weekend vs. weekday and home vs. work), staying within the same range from 0 (when there is only one profile) to 1 (when there are multiple profiles, and all are equally important).
4.6 HoWDe parameters details
$\mathbf{dhn}$: First, to account for a minimum level of information within a day, we focus only on days with hour sequences consisting of at least three stops, i.e. stops visited in at least three hours of the day (i.e. days with at least three hours of data). The HoWDe algorithm, however, can be tuned to handle different thresholds based on the specific needs or constraints in place.
$\mathbf{dn_{H,W}}$: Second, to account for the potentially varying data quality across different datasets, we introduce two independently configurable pairs of parameters for home and work locations. The first pair of parameters works as a maximum fraction of days without data per sliding window, marking a day with an “undetected home/work location” if the fraction of days without data within the sliding window exceeds the threshold.
$\mathbf{hf_{H,W}}$: A third pair of parameters threshold the minimum fraction of hours per night or day period for a location to be home or work respectively. If the average fraction of visits across all days in the sliding window is below the threshold, the location cannot be considered a potential home/work.
$\mathbf{df_{H,W}}$: Fourth, while home locations are expected to be visited almost every day (provided that enough data are available), work locations are likely to follow a different pattern, as those locations will be subject to local work habits, e.g. four days working week. To account for regularities across days, we introduce an additional requirement for work locations: a work location should be visited for a minimum fraction of days with data.
Table 2: Summary description of the HoWDe configurable parameters. The table reports the parameter names, whether they are used in home detection, work detection, or both, and the description of the parameter purpose. When looking at the hourly bins with data (hours with data), that is the case of $dhn$, $hf_{H}$, and $hf_{W}$, counts and fractions are computed with respect to night/business hourly bins only. Note that, $dhn$, $dn_{H}$, and $dn_{W}$ act as minimum activity thresholds for a given day to have an estimated home or work location estimated. $hf_{H}$, $hf_{W}$, and $df_{W}$ serve a similar purpose but for location-specific visit requirements allowing each location above the threshold to be a candidate home/work location. Specifically, fractions of visits to home/work location candidates are computed relative to the total hours in which an individual has recorded data during the specified period.
Name Usage Description
$\mathbf{\Delta_{T}}$ Home/Work data quality Size of the memory window
$\mathbf{dhn}$ Home/Work data quality Min. number of night/business hourly-bins with data in a day
$\mathbf{dn_{H}}$ Home data quality Max. fraction of days without data in a window
$\mathbf{dn_{W}}$ Work data quality Max. fraction of days without data in a window
$\mathbf{hf_{H}}$ Home estimation Min. fraction of night hourly-bins (avg. over days in the window)
$\mathbf{hf_{W}}$ Work estimation Min. fraction of business hourly-bins (avg. over days in the window)
$\mathbf{df_{W}}$ Work estimation Min. fraction of days within the window
For details on parameter implementation, we refer to the Sec. SI-6, where we provide the algorithm’s pseudocode.
4.7 Accuracy and “fraction of not detected” definition
In this work, we report the performance of the algorithm in terms of two main quantities: “accuracy” and “fraction of not detected”. Here, we provide the definitions of these two concepts. The definitions below are explained in the case of the home detection, and they apply in the same way for the work detection.
The fraction of ‘not detected’ indicates cases where a home location was labelled, but the algorithm couldn’t assign a home due to the thresholds on data quality or minimum fraction of visits. In the case of D1, the dataset consists of a set of user-weeks, corresponding to users that self-reported living in a specific home (the “ground truth home”) in a given week. For some of these user-weeks, our algorithm identifies a “detected home”. We evaluate the performance of the algorithm by comparing pairs of “ground truth home” and “detected home” across user-weeks. The “fraction of not detected” is thus computed as $f_{ND}=1-n_{\bar{H}}/n_{H}$, where $n_{\bar{H}}$ is the number of user-week pairs with a detected home, and $n_{H}$ is the number of user-week pairs with a ground truth home. In the case of D2, where ground truth information is not referring to a specific week, $n_{\bar{H}}$ corresponds to the number of users with a detected home, and $n_{H}$ corresponds to the number of users with a ground truth home.
The “accuracy” captures the fraction of correctly labelled home locations among the set of detected home locations. Therefore, we define accuracy as $Acc=m_{\bar{H}}/n_{\bar{H}}$, where $m_{\bar{H}}$ is the number of user-weeks (for D1) or users (for D2) for which the detected home location matches the ground truth home location. In dataset D2, if annotators identified multiple homes for an individual, we considered it a match if the individual’s detected home location corresponded to at least one of the annotated ground truth home locations. Note that our definition does not correspond exactly to the traditional definition of accuracy because we separately assess the algorithm performance in terms of the detection/non-detection of a home (captured by $f_{ND}$) and the detection of the correct home (captured by $Acc$).
The definitions above are also used in the case of work detection, where only users with a reported work location, $n_{W}$, are included in the denominator of $f_{ND}$, and only those with a detected and a matching work location are counted in the nominator and in the accuracy computation, as $n_{\bar{W}}$ and $m_{\bar{W}}$, respectively.
5 Data and Code Availability
Data to ensure replicability of the analysis performed in this work can be found at
https://doi.org/10.11583/DTU.28846325 [Note: The link will become available upon publication]. To protect user anonymity, we applied the following privacy-preserving transformations to the stop sequences. First, time was discretized into 10-minute intervals. Second, each user’s timeline was shifted so that their first recorded activity occurred in the first week of 1970. Third, entire days were randomly shuffled within each week, while preserving the distinction between weekdays and weekends. All code needed to implement HoWDe is available at [37]. We also provide HoWDe as a Python (PyPI) package at [54].
Acknowledgments
L.L. has been supported by the ERC project “IMMUNE” (Grant agreement ID: 101003183). L.L. acknowledges the support from the “Fondazione Romeo ed Enrica Invernizzi” for the research activities of the ‘Covid Crisis Lab’ at Bocconi University. L.A. and S.D.S. acknowledge support from the Independent Research Fund Denmark (DFF) through the DFF-Research Project 1 (Inge Lehmann) grant “Gender Gaps in Human Mobility”. The authors thank Y. Liao for providing thoughtful feedback on the code release, which helped improve its clarity and usability.
Declaration of Interest statement
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
Author Contributions
S.D.S., L.L., S.F., and L.A. conceived the study. S.D.S., L.L., and O.D.L.C. curated and preprocessed the mobility and ground truth data. Two alternative algorithms were independently developed by L.L., S.F., and O.D.L.C., and by S.D.S. and L.A. The two designs were compared and discussed collaboratively, and the final HoWDe workflow is based on the design by S.D.S. and L.A., with L.L. contributing to its finalization. S.D.S. and L.L. designed the analysis pipeline and conducted the statistical analyses. S.F. and L.A. provided domain expertise and guidance on the interpretation of results. S.D.S. and L.L. drafted the manuscript. All authors discussed the results and approved the final version of the manuscript.

### v2

@@
-Establishing validated standards for Home and Work location Detection
+HoWDe: a validated algorithm for Home and Work location Detection
 Abstract
-Smartphone location data have transformed urban mobility research, providing unprecedented insights into how people navigate and interact in cities.
-However, leveraging location data at scale presents methodological challenges.
-Accurately identifying individuals’ home and work locations is critical for a range of applications, including commuting analysis, unemployment estimation, and urban accessibility studies.
-Despite their widespread use, home–work detection methods lack a standardized framework that accounts for differing data quality and that is validated against ground-truth observations.
-This limits the comparability and reproducibility of results across studies and datasets.
-In this paper, we present HoWDe, a robust algorithm for identifying home and work locations from mobility data, explicitly designed to handle missing data and varying data quality across individuals.
-Using two unique ground-truth datasets comprising over $5,100$ individuals from more than $80$ countries, HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across countries and demographic groups.
-We examine how parameter choices shape the trade-off between accuracy and user retention, and demonstrate how these methodological decisions influence downstream applications such as employment estimation and commuting pattern analysis.
-By supporting in-house pre-processing through a transparent and validated pipeline, HoWDe also facilitates the sharing of privacy-preserving mobility data.
-Together, our tools and findings establish methodological standards that support more robust, scalable, and reproducible mobility research at both individual and urban scales.
-Keywords: GPS data, Home and work detection, Human Mobility, Stop-location sequences, Mobility algorithms, Missing data, Urban analytics
+Smartphone location data have become a key resource for understanding urban mobility, yet extracting actionable insights requires robust and reproducible preprocessing pipelines.
+A central step is the identification of individuals’ home and work locations, which underpins analyses of commuting, employment, accessibility, and socioeconomic patterns.
+However, existing approaches are often ad hoc, data-specific, and difficult to reproduce, limiting comparability across studies and datasets.
+We introduce HoWDe, an open-source software library for detecting home and work locations from large-scale mobility data.
+HoWDe implements a transparent, modular pipeline explicitly designed to handle missing data, heterogeneous sampling rates, and differences in data sparsity across individuals.
+The code allows users to tune a small set of interpretable parameters, enabling to adapt the algorithm to diverse applications and datasets.
+Using two unique ground truth datasets comprising 5,0995{,}099 individuals across $68$ countries, we show that HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across demographic groups and geographic contexts.
+We further demonstrate how parameter settings propagate to downstream metrics such as employment estimates and commuting flows, highlighting the importance of transparent methodological choices.
+By providing a validated, documented, and easily deployable pipeline, HoWDe supports scalable in-house preprocessing and facilitates the sharing of privacy-preserving mobility datasets.
+Our software and evaluation benchmarks establish methodological standards that enhance the robustness and reproducibility of human mobility research at urban and national scales.
+Keywords: Home and work detection, Open-source, Human Mobility, GPS data, Stop-location data, Missing data, Urban analytics
 1 Introduction
-Location data collected from smartphones has become indispensable for human mobility analyses [1, 2].
-These datasets allow studying large samples of individuals globally, providing insights on fundamental aspects of human movements [3], epidemic spreading [4, 5, 6, 7], migration patterns [8, 5], transportation modeling [9], urban planning [10, 11], and the impact of natural disasters and societal shocks on population movements [12, 13, 14, 15, 16].
-In light of these advancements, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [17].
-A critical component of location data analyses involves the detection of individuals’ visits to their residential [18, 19, 20], and work location [17].
+Location data collected from smartphones has become a key resource for the analysis of urban systems and human mobility [1, 2, 3, 4, 5, 6].
+These datasets enable the study of large, diverse populations and have been used within a range of areas, including the study of human movements [7, 8, 9], epidemic spreading [10, 11, 12, 13], migration patterns [14, 11], transportation modeling [15], urban planning [16, 4, 5, 6], and assessing the impact of natural disasters and societal shocks on population movements [17, 18, 19, 20, 8].
+As the availability of such data increases, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [21].
+A critical component of location data analyses is the detection of individuals’ residential [22, 23, 24], and work locations [21].
 The accurate identification of home and work locations is key for mobility research, as it enables the analysis of factors such as socio-economic background [25, 26, 20], commuting behaviors [27, 8], social contact networks [28, 19], and accessibility to amenities and services [29, 19].
 More broadly, it contributes to advancements in computational demography and social science research [30, 21].
-The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behavior, such as the tendency to spend nights at home and working hours at a workplace [27, 28, 29].
-These methods face significant challenges.
-First, they are sensitive to missing data, which is common in smartphone records [17].
-Data quality varies across individuals and over time, as location gaps arise due to factors like inactivity, disabled services, or poor signal, introducing biases in home and work inference [30, 31, 17].
-Second, they rely on the choice of parameters, such as time span required to establish a home location or the minimum number of hours per day spent at work.
-Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains unclear, and there is no consensus on how to choose one specific algorithm and its parameters’ values [21, 32, 33, 34, 35, 17], in part due to a lack of ground truth data.
-Validation efforts so far have been limited, focusing on small subsets of individuals, home detection alone, and datasets with low spatial resolution and short time periods [34, 35, 32, 21].
-Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been understood in relation to algorithmic performance.
-The lack of thorough validation has hindered the refinement and harmonization of these methods into standardized best practices [36, 17].
-In turn, the absence of such standards can impact the comparability and reproducibility of mobility studies [17].
-In this work, we leverage two distinct ground truth datasets to address these key challenges in home/work location detection.
-Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 (see Methods for details).
-First, we leverage these data to uncover behavioral patterns in home and work visits.
-This allows us to empirically assess whether the assumptions underlying home and work detection hold in practice.
-Additionally, it enables us to quantify the variability in home-work visit behavior across a diverse set of users.
-Second, we formalize HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach.
+The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behaviour, such as the tendency to spend nights at home and working hours at a workplace [31, 32, 33].
+However, these methods face significant challenges.
+First, they are sensitive to missing data, which often stems from inactivity or poor signal and can bias home and work detection [34, 35, 21].
+Second, they rely on the choice of parameters, such as the time span required to establish a home location or the minimum number of visits required to identify a valid work location.
+These limitations underscore the need for methods that systematically address both the heterogeneity of missing data within individuals and the sensitivity to heuristic parameter definitions [36, 21].
+Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains insufficiently validated, and there is no consensus on how to choose one specific algorithm and its parameters’ values [25, 37, 38, 39, 40, 21], in part due to a scarcity of ground truth data.
+Existing validation efforts have been limited to small samples, short time periods, low spatial resolution, or to home detection alone [39, 41, 40, 37, 25].
+Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been studied systematically.
+Limited validation has hindered the refinement and harmonisation of these methods into standardised best practices  [21].
+In turn, the absence of standard algorithms hinders the comparability and reproducibility of mobility studies  [42, 21].
+In this study, we leverage two distinct ground truth datasets collected across more than 68 countries to address these key challenges in home/work location detection.
+Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 and 2021.
+First, we formalise HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach.
 HoWDe combines the explicit detection of both home and work locations [19, 33], the identification of the workplace based on recurrent visits during working hours [33], the identification of the home as the most frequently visited location at night [40, 32, 39, 25], a sliding window to identify home and job changes over time [19], with tunable window length [19].
-Additionally, HoWDe improves upon existing approaches by better accounting for data gaps and their variation across users and over time.
-We provide an open-source implementation of the algorithm [37].
-Finally, we assess HoWDe’s performance against the ground truth data.
-We provide a detailed analysis of the role of its parameters, and we study how performance changes across age groups, gender, countries, and urbanization levels.
-We conclude by quantifying how errors in the estimation of home and work locations propagate to key downstream tasks such as estimating employment rates or commuting distances.
-2 Results
-2.1 Unveiling temporal patterns of home-work visits
-We analyze the regularity of home and work visits by measuring how often and when individuals are recorded at their annotated locations.
+Additionally, HoWDe improves upon existing approaches by accommodating users whose data sparsity varies over time.
+This is achieved by identifying home and work locations based on the proportion of time spent at a given location, while controlling for individual-level sparsity rather than excluding users based on absolute time or visit thresholds, as in prior work.
+Second, we assess HoWDe’s performance against the ground truth data.
+We provide a detailed analysis of the role of its parameters, and we study how performance changes across age groups, gender, countries, and urbanisation levels.
+We quantify how errors in the estimation of home and work locations propagate to key downstream tasks such as estimating employment rates or commuting distances.
+The article is organized as follows: in Section 2 we introduce existing approaches to home and work location detection; in Section 3 we introduce the ground truth data and additional methods used in this study; in Section 4 we explore the ground truth data to motivate the heuristic choices of HoWDe, in Section 5 we describe the HoWDe algorithm; in Section 6 we evaluate the performance of the algorithm, explore the effect of parameter choices, assess robustness across demographic groups, and demonstrate applications in estimating employment rates and commuting patterns.
+2 Background
+Smartphones use an in-built GPS sensor to estimate the device’s latitudes and longitude coordinates over time, at a frequency that varies depending on device usage.
+The device position is then collected by various location-based apps, such as navigation or social media apps.
+These data can be used by location intelligence companies or researchers to study human mobility and its regularities, both at the individual and at the population level.
+Raw location data utilised in urban and mobility studies consists of timestamped sequences of (latitude, longitude) records associated with a unique user ID.
+The first pre-processing step involved in the analysis is the so-called “stop-location detection" [43, 44].
+Stop location detection consists of (i) identifying records that are captured when individuals stay within a small area for a sufficient amount of time (e.g.
+when they are at home, at the office, or in a restaurant); (ii) applying spatial clustering to all these “static records", to identify a set of unique locations, i.e.
+those places where people spend time.
+Stop location definition might vary depending on the type of data analyzed and the specific characteristics of the data points, e.g., whether you use Call Detail Records (CDR), eXtended Detail Records (XDR), or GPS points.
+However, the principled choices these algorithms are founded on can naturally be extended to any of these different cases, making them easily generalizable.
+In this work, stop locations are defined as spatiotemporally close sets of GPS points.
+For further details, we refer the reader to the Supplementary material (Section S-2) and prior work [43, 19].
+After detecting the individuals’ stop locations, individual trajectories can be summarised as sequences of stays.
+Each stay is associated with a start time (when the stay started), a duration (how long the stay lasted), and a location ID (where the stay happened).
+A common next step in many mobility studies involves identifying key locations, such as individuals’ homes and workplaces.
+To achieve this, researchers and practitioners typically assume that visits to home and work locations exhibit specific daily and weekly regularities.
+There are various approaches to identifying home and work locations.
+The majority of studies adopt ad-hoc heuristics [37, 25, 38, 40, 33, 45, 46].
+The most commonly used heuristics rely on rules in relation to “how much”, and “when” an individual visits each location during the period of observation [37, 25, 38, 47, 48, 49].
+Some studies complements temporal heuristics with semantic information, for example by matching stay locations to nearby points of interest (POIs) or land-use categories to ensure that inferred homes fall within residential areas and work locations within commercial or industrial zones [50, 51].
+Importantly, in most cases, these heuristics are not validated against ground truth data nor are standardised across studies.
+A minority of studies, instead, leverage machine learning techniques [52, 41]; however, these methods either generalise poorly to individuals not seen during training [52], or have been tested only in specific cities and rely on a particular mode of transport, such as private cars [41].
+In this work, we compare our algorithm against two established baseline methods commonly used in mobility research, selected for their widespread adoption and their relevance to circadian and regularity-based home-work detection heuristics.
+These methods were designed and utilised over the past decade, leveraging diverse sets of mobile-phone-based data (e.g., Call Detail Records (CDRs), eXtended Detail Records, (XDRs), Control Plane Records (CPRs), or GPS records) [25, 53, 54, 40] as well as activity-based travel surveys [42].
+Among the home/work detection approaches in the literature, these are the most widely used with GPS data because they rely solely on circadian regularities, making them easy to deploy across heterogeneous mobility datasets without requiring POIs, land-use information, or detailed ground truth.
+Both approaches build on the rationale of ranking locations based on the day of the week (weekdays vs.
+weekends) and the time of day (daytime vs.
+nighttime) when visits occur.
+In the following, we describe these two methods in detail.
+2.1 The Atlas approach
+The method used in [3, 32, 55] (hereinafter Atlas method) defines a home location as the most probable location visited between 10 p.m.
+and 6 a.m., i.e., the stop location where a device was found spending most of the nighttime.
+The scikit-mobility Python library—a standard toolkit for mobility analysis—implements this definition [56].
+To ensure that only users with sufficient data quality are included in the analysis, a filtering strategy is typically applied.
+Individuals who spend fewer than 10 nights in the same census block group during a six-month observation period [3] (or fewer than 5 nights over a three-month period [55]) are excluded from the analysis.
+The Atlas method does not explicitly estimate work locations; however, its approach to identifying home locations can be extended to infer work.
+Following a similar rationale, we define an Atlas work location as the most common location on weekdays between 9 a.m.
+and 4 p.m.
+2.2 The TimeGeo approach
+More complex heuristics are adopted in [33] (hereinafter referred to as the TimeGeo method), in particular for identifying work locations.
+TimeGeo ranks stop locations based on the frequency of visits during weekday nights (between 7 p.m.
+and 8 a.m.) and weekends and picks the most visited location as an individual’s home.
+The same logic is applied when detecting the work location, but now focusing on weekday work time periods (between 8 a.m.
+and 7 p.m.) and including additional requirements: i) the candidate work location is visited at least three times during the study period, and ii) the distance between the candidate work location and the individual’s home location exceeds 500 meters.
+Similar to the Atlas method, TimGeo applies a data quality filter, excluding users with fewer than 50 total stops and fewer than 10 home stays during the observation period.
+Atlas and TimeGeo rely on similar strategies for handling missing data.
+After estimating home and work locations, users are filtered based on their overall activity levels and/or the number of nights spent at home.
+While this filtering excludes individuals with low mobility or no stable residence, it has key limitations.
+First, existing methods fail to account for temporal heterogeneity in data availability within individuals.
+Gaps in location data—caused by device inactivity, disabled services, or poor signal—are common in smartphone records and vary widely across users and within users over time  [57, 35, 21].
+These inconsistencies can introduce systematic biases in the inference of home and work locations, particularly when left unaccounted for [21].
+Second, it relies on parameter choices that are typically ad hoc, lacking empirical grounding, and are often based on visit counts, a design that can disproportionately impact users with shorter or less consistent observation periods.
+Moreover, systematic large-scale validation using GPS data remains limited [21].
+Most existing evaluations rely on lower-resolution sources such as CDRs, XDRs, and CPRs [40], which constrain temporal precision and spatial accuracy.
+HoWDe addresses these gaps by grounding its design in large-scale GPS data annotated with home and work locations.
+The methodology we propose integrates established processing techniques while explicitly addressing the challenges of data sparsity and parameter sensitivity, offering a standardised and scalable approach to home–work location detection.
+3 Materials and Methods
+3.1 Ground truth datasets
+To validate the proposed methods, we use two GPS-based datasets with ground truth home and work labels: one featuring large-scale self-reported annotations, and the other comprising expert-labelled trajectories.
+Together, they enable robust evaluation across diverse populations and settings.
+3.1.1 Dataset 1: DTU stop data and self-reported annotations
+This dataset uses mobility data collected between 2017 and 2019, previously analyzed in [7], where stop locations were identified using the scalable stop-location detection algorithm InfoStop [43].
+The stop location data captures the mobility patterns of 96,065 users aged 25 to 64 years old.
+The dataset includes $2,309$ self-reported home and $2,479$ work location labels, collected from $4,599$ mobile phone and smart band users from 68 countries across continents.
+The labelled sample is biased towards urban areas and higher- to middle-income countries.
+To preserve privacy, spatial coordinates have been removed (See Section 8 for data availability).
+3.1.2 Dataset 2: Veraset Movements data and annotation
+This dataset is based on anonymised GPS location data provided by Veraset as part of the World Bank’s “Monitoring COVID-19 Policy Response through Human Mobility Data” project.
+Unlike most GPS providers, Veraset sources its data from thousands of Software Development Kits (SDKs), reducing sampling bias and covering approximately $5\%$ of the global population [58].
+The data spans 24 months, from January 1, 2020, to December 31, 2021, and has been previously used to assess the impact of the pandemic and the policy responses [20].
+Dataset D2 includes GPS trajectories from $100$ devices in each of five countries: Brazil, Colombia, Indonesia, Mexico, and the Philippines.
+For each country, users were selected to reflect high-, medium-, and low-wealth neighbourhoods ($33\%$), based on a wealth classification framework developed for GPS mobility analysis in middle-income countries [20].
+To ensure data quality, users were required to have stop records on at least 20% of days both before the WHO’s pandemic declaration and over the 2020 pandemic period.
+For each user, experts annotated the stop location sequences for 2020 by inspecting location visits, satellite imagery, and local amenities (see Supplementary material Section S-1 for a detailed description of the annotation procedure).
+Throughout the analysis, we extended these annotations to 2021 data to increase the amount of available location data for the selected users.
+We note that data for 2021 contains substantially fewer observations and exhibits marked shifts in data quality associated with pandemic-related changes in mobility behaviour.
+The final sample includes 500 users, of whom $287$ have at least one work location labelled.
+Since annotators inspected the complete longitudinal stop sequence, it is not possible to explicitly consider changes in home and work for each individual.
+3.2 Measuring ground truth home/work visits to inform HoWDe design
+In Section 4, we examine empirical regularities in home and work visits using the ground truth data.
+Here, we detail the methods used to analyse these patterns.
+First, we outline how we identified typical daily patterns of home and work visits, which we refer to as behavioural clusters.
+Second, we describe how we assessed whether individuals in our dataset are captured by a single behavioural cluster or distributed across multiple clusters.
+3.2.1 Home and work visits behavioural clusters
+To characterise daily regularities in home and work visits, we represent each user’s behaviour as a sequence of 24 hourly bins, corresponding to one day.
+Each hourly bin is assigned one of three categories: (1) visit to the home (or work) location, (2) a visit to another location (e.g., work, leisure, or other activities), and (3) missing data.
+For home patterns, category (1) corresponds to being at the home location; for work patterns, it corresponds to being at the workplace, while hours spent at home are labelled as visits to another location (2).
+These daily sequences capture the temporal structure of individuals’ visits, allowing us to quantify variation in visit regularity across users and days.
+We cluster these sequences using the K-Modes algorithm—a categorical adaptation of K-Means [59]—to identify days exhibiting similar visitation patterns, which we refer to as behavioural clusters.
+The number of clusters is determined using the elbow method.
+Rather than defining a strict typology, here our aim is to uncover dominant behavioural profiles that summarise the most common home and work visitation patterns (see Section 4).
+These clusters form the basis for assessing whether individuals’ behaviour aligns with a single dominant pattern or spans multiple profiles, informing the design of HoWDe.
+3.2.2 Behavioral profile entropy
+To quantify how consistently users follow specific behavioural profiles, we measure the normalised Shannon entropy [60], treating each behavioural cluster as a distinct state.
+The entropy is defined as:
+$H_{\text{norm}}=-\frac{1}{\log{k}}\sum_{i=1}^{k}p_{i}\log{p_{i}},$ (1)
+where $k$ is the total number of possible clusters, and $p_{i}=\frac{n_{i}}{N}$ denotes the probability of a user being assigned to cluster $i$, with $n_{i}$ being the number of days classified in that cluster, and $N$ the total number of days for the user.
+Normalising by the maximum entropy $\frac{1}{\log{k}}$ ensures comparability across different settings—such as weekdays vs.
+weekends, or home vs.
+work patterns, by bounding the entropy between 0 (only one profile used) to 1 (maximum diversity; multiple profiles and all used equally).
+3.3 Performance evaluation metrics
+In this work, we report the performance of the algorithm in terms of: “detected accuracy” and “fraction of not detected”.
+Below, we define these metrics in the context of home detection; the same definitions apply analogously to work detection.
+The fraction of ‘not detected’ indicates cases where a home location was labelled, but the algorithm couldn’t assign a home due to thresholds on data quality or a minimum fraction of visits.
+In the case of D1, the dataset consists of a set of user-weeks, corresponding to users who self-reported living in a specific home (the “ground truth home") in a given week.
+For some of these user-weeks, our algorithm identifies a “detected home".
+We evaluate the performance of the algorithm by comparing pairs of “ground truth home" and “detected home" across user-weeks.
+The “fraction of not detected” is thus computed as $f_{ND}=1-n_{\bar{H}}/n_{H}$, where $n_{\bar{H}}$ is the number of user-week pairs with a detected home, and $n_{H}$ is the number of user-week pairs with a ground truth home.
+In the case of D2, where ground truth information does not refer to a specific week, $n_{\bar{H}}$ corresponds to the number of users with a detected home, and $n_{H}$ corresponds to the number of users with a ground truth home.
+In state-of-the-art methods, such as Atlas and TimeGeo, home locations are always assigned to users who are included in the analysis, resulting in no undetected cases.
+However, this practice can introduce systematic errors when data coverage biases are present throughout the study period, e.g., due to data-provider issues, external shock events, or individual changes in activity levels.
+The “detected accuracy” captures the fraction of correctly labelled home locations among the set of detected home locations, following the approach previously applied by [40].
+Therefore, we define detected accuracy as $Acc=m_{\bar{H}}/n_{\bar{H}}$, where $m_{\bar{H}}$ is the number of user-weeks (for D1) or users (for D2) for which the detected home location matches the ground truth home location.
+In dataset D2, if annotators identified multiple homes for an individual, we considered it a match if the individual’s detected home location corresponded to at least one of the annotated ground truth home locations.
+Note that our definition does not correspond exactly to the traditional definition of accuracy because we separately assess the algorithm performance in terms of the detection/non-detection of a home (captured by $f_{ND}$) and the detection of the correct home (captured by $Acc$).
+The definitions above are also used in the case of work detection, where only users with a reported work location, $n_{W}$, are included in the denominator of $f_{ND}$, and only those with a detected and a matching work location are counted in the nominator and in the detected accuracy computation, as $n_{\bar{W}}$ and $m_{\bar{W}}$, respectively.
+3.4 Comparing home/work detection methodologies
+We compare the performance of three home/work detection methods, HoWDe, Atlas, and TimeGeo (see Section 2), within a shared evaluation framework.
+All methods are applied to an identical set of stop locations, generated using the same processing pipeline for each dataset (see Supplementary material, Section S-2).
+As described in Section 2, both Atlas and TimeGeo typically apply pre-filtering steps before location detection.
+Although HoWDe includes an internal data-quality assessment, we pre-filter users for all methods to ensure comparability.
+Specifically, we retain only users with at least 10 days of data in dataset D1 (note that $83\%$ of these users have $\geq 60$ days) and at least 71 days in dataset D2 (applicable to all D2 users).
+This filtering step ensures that performance differences reflect methodological rather than data-quality variations.
+To further isolate methodological differences from temporal definitions, we adopt consistent nighttime (home) and daytime (work) windows across all methods, preserving the weekend–weekday distinctions defined in the baselines (see Section 2).
+These time ranges are empirically determined using ground truth data (see Section 4).
+4 Empirical regularities in home and work visits
+To establish validated standards for home and work location detection, we first examine the temporal patterns that characterise visits to home and work locations in our ground truth dataset.
+The HoWDe algorithm (presented in Section 5 below) is built on these insights.
+We analyse the regularity of home and work visits by measuring how often and when individuals are recorded at their annotated location (see Section 3.1).
 Figure 1a-b shows the probability of being at home or work by hour of the day.
-We find that home visits peak between midnight and 6 AM, while work visits are most frequent during typical business hours (between 9 AM and 4 PM).
-Patterns of visits to the home locations are similar between weekdays and weekends (see Figure 1c-e).
-For the work location, we observe a drastic reduction in the fraction of individual visiting their work location during weekends (see Figure 1d-f).
-Results are consistent across both datasets, with overall higher probabilities of being at home (and lower probabilities of being at work) found in D2 due to the COVID-19 pandemic containment policies in place during most of the period covered by the data (see Sec.SI-3).
+We find that home visits peak between midnight and 6 a.m., while work visits are most frequent during typical business hours (between 9 a.m.
+and 4 p.m.).
+Patterns of visits to the home locations are similar between weekdays and weekends (see Figure 1a).
+For the work location, we observe a drastic reduction in the fraction of individual visiting their work location during weekends (see Figure 1b).
+Results are consistent across both datasets, with overall higher probabilities of being at home (and lower probabilities of being at work) found in D2 due to the COVID-19 pandemic containment policies in place during most of the period covered by the data (see Supplementary material Sec.
+S-3).
 Our analysis suggests that work location is best determined using weekday daytime hours, while home location can be identified using night hours, regardless of weekdays or weekends, similar to previous approaches [3, 32, 40].
-However, while general trends exist, we also find significant variations across days and users, highlighting pronounced behavioral heterogeneities at the individual level.
-Figure 1c-f reports disaggregated visit patterns for single “user-day” pairs.
-Most users’ daily patterns can be grouped into six behavioral profiles based on home and three on workday visits (see Methods for details on how behavioral profiles are obtained).
+However, while general trends exist, we also find significant variations across days and users, highlighting pronounced behavioural heterogeneities at the individual level (Figure 1a-b).
+Figure 1 panels c-f report disaggregated visit patterns for “user-day” pairs, where each vertical line represents one person’s hourly sequence of visits to the home or work location over a single day.
+Most individuals’ daily patterns can be grouped into distinct behavioural profiles.
+We identify six profiles based on home visits and three based on workday visits (see Section 3.2.1 for details on how behavioural profiles are obtained).
 Figure 1: Behavioral visit profiles to home and work locations a.
@@
 The user-day profiles are sorted by cluster, hour of the first home visit, total time spent at home, and hours without data.
-Clusters are separated by black vertical lines, and the fraction of days per cluster is noted on the x-axis.
-With respect to home visits, most user days fall into one of three typical profiles: i) days when users mostly stay at home, ii) when they are rarely/never at home, and iii) when they are away from home during typical business hours (see Figure 1c,e).
-Similarly, work-behavior profiles are described by three main daily patterns: i) days when users spend business hours at work, ii) when they are rarely/never at work, and iii) when they spend night hours at work (Figure 1d,f).
-We find that the distribution of individuals across these profiles is generally independent of their socio-demographic characteristics, such as age, gender, and urbanization (see Sec.SI-5).
-As the behavioral profiles we identified are based on single user-day activity data, an important question is whether individuals switch between these profiles or, rather, a single profile consistently represents them.
+Clusters are separated by black vertical lines, labelled with roman numerals, and the fraction of days per cluster is noted on the x-axis.
+With respect to home visits, three profiles account for most user-days: i) days when users mostly stay at home, ii) when they are rarely/never at home, and iii) when they are away from home during typical business hours (see Figure 1c,e).
+Similarly, work-behaviour profiles are described by three main daily patterns: i) days when users spend business hours at work, ii) when they are rarely/never at work, and iii) when they spend night hours at work (Figure 1d,f).
+We find that the distribution of individuals across these profiles is generally independent of their socio-demographic characteristics, such as age, gender, and urbanisation (see Supplementary material Sec.S-4).
+As the behavioural profiles we identified are based on single user-day activity data, an important question is whether individuals switch between these profiles or, rather, a single profile consistently represents them.
 We find that users visit a median of $5$ out of $6$ home clusters and $2$ out of $3$ work clusters (for both D1 and D2).
-We measure the individual inter-profile variability through the normalized entropy, a measure ranging from 00 (when all days of a user fall into a single profile) to $1$ (when all profiles are equally important) [38] (see Methods).
-On average, users exhibit a normalized entropy of $0.66\pm 0.01$ for home-profile changes between weekdays and $0.48\pm 0.01$ for weekends ($0.62\pm 0.01$ and $0.42\pm 0.01$ for D2 respectively), and $0.51\pm 0.01$ for work-profile changes on weekdays only ($0.40\pm 0.02$ for D2).
-These values are comparable to the ones obtained when randomly allocating profiles among user days, confirming that users do not have a single dominant behavioral profile (see Table si.1).
-All in all, our findings suggest that focusing on the same profiles across users is unlikely to favour one specific socio-demographic group, thus supporting the choice of focusing on weekday daytime hours (9 am to 4 pm) for detecting work locations and nighttime hours (12 am to 6 am) for detecting home locations, regardless of weekdays and weekends.
+We measure the individual inter-profile variability through the normalised entropy, a measure ranging from 0 (when all days of a user fall into a single profile) to $1$ (when all profiles are equally important) [60] (see Sec.
+3.2.2).
+On average, users exhibit a normalised entropy of $0.66\pm 0.01$ for home-profile changes between weekdays and $0.48\pm 0.01$ for weekends ($0.62\pm 0.01$ and $0.42\pm 0.01$ for D2, respectively), and $0.51\pm 0.01$ for work-profile changes on weekdays only ($0.40\pm 0.02$ for D2).
+These values are comparable to the ones obtained when randomly allocating profiles among user days, confirming that users do not have a single dominant behavioural profile (see Table S-1, in Supplementary material Sec.
+S-5)..
+All in all, our findings suggest that focusing on the same profiles across users is unlikely to favour one specific socio-demographic group, thus supporting the choice of focusing on weekday daytime hours (9 a.m.
+to 4 p.m.) for detecting work locations and nighttime hours (12 a.m.
+to 6 a.m.) for detecting home locations, regardless of weekdays and weekends.
 These time windows, combined with minimum data coverage requirements, form the foundation of the heuristics implemented by the HoWDe algorithm.
-2.2 HoWDe: a Home and Work Detection algorithm
-In this section, we introduce HoWDe, a dynamic home and work detection algorithm that can capture residential moves and job changes over time through a sliding window approach.
-A key feature of HoWDe is the identification of home and work locations based on the proportion of time spent at a given location, rather than the absolute amount of time as approached in previous literature [29, 28, 14, 15, 33, 34].
-This enables HoWDe to accommodate users with varying amounts of missing data over time (see Figure 1c-f), through tunable parameters that set data-quality filters.
-We share an open-source implementation of the algorithm in PySpark [37].
-2.2.1 HoWDe workflow
+5 HoWDe: a Home and Work Detection algorithm
+In this work, we introduce HoWDe, a dynamic home and work detection algorithm that captures residential moves and job changes over time using a sliding window approach.
+The heuristics adopted by HoWDe build on the empirical results obtained in the previous Section 4, leveraging our comprehensive ground truth data.
+A key feature of HoWDe is its ability to handle data sparsity through a series of internal steps designed to ensure robust home and work detection even when individual traces are incomplete.
+Rather than relying on absolute time spent at a location [33, 32, 19, 20, 38, 39], HoWDe identifies locations based on the proportion of time spent, a choice that naturally accommodates users with uneven data sparsity over time (see Section 4).
+This proportional representation is combined with fixed-length hourly sequences and tunable temporal coverage filters, allowing the algorithm to standardize temporal structure across days while explicitly accounting for missing data.
+Together, these design choices enable HoWDe to remain robust under varying levels of sparsity both across and within individuals.
+We provide an open-source implementation of HoWDe in PySpark [61].
+Supplementary materials section.
+S-6 outlines all aspects of the software implementation, including availability, packaging, licensing, engineering decisions, and the development model.
+We further emphasise computational efficiency in Sec.
+S-6.4, where we compare the resource utilisation of HoWDe with that of the benchmark methodologies (Atlas and TimeGeo).
+The next subsections detail the algorithm workflow (Section 5.1) and its tunable parameters (Section 5.2).
+5.1 HoWDe workflow overview
 HoWDe takes as input a user’s pre-processed sequence of stops, where each stop is fully defined by a user-location unique ID and a start and end timestamp.
 First, we perform a temporal aggregation of the stop sequences into hourly-bin sequences.
-In each hourly bin, we retain the location where the individual has spent most of the time.
-This aggregation limits sequences to a maximum of 24 locations per day, boosting computational efficiency and making HoWDe faster than previous approaches based on sliding windows, with only limited information loss [14, 15].
-Second, we exclude days with insufficient hourly bins with data (this step is controlled by the tunable parameter $dhn$, see Table 2).
+In each hourly bin, we retain the location where the individual spent the most time.
+This aggregation limits sequences to a maximum of 24 locations per day, boosting computational efficiency and making HoWDe faster than previous approaches based on sliding windows [19, 20], with only limited information loss.
+An extended evaluation of how increasing temporal resolution beyond 1-hour bins affects performance is provided in Supplementary materials sec.
+S-7.
+Second, we exclude days with insufficient hourly bins with data (this step is controlled by the tunable parameter $C_{hours}$, see Table 1).
+This step accounts for the observation that the amount of missing data varies widely across users and over time, and even within individual users’ routines (see Section 4).
 This is done separately for work detection (where we consider business hours) and home detection (where we consider night hours, see Figure 2b).
-Hence, it is possible that one day is retained only for the home detection or for the work detection.
-Third, for each day $t$, we calculate the fraction of hourly-bins spent at each location during nighttime (24 pm to 6 am for home detection), and typical business hours (9 am to 4 pm on weekdays, for work detection).
+Hence, it is possible that one day is retained only for home detection or for work detection.
+Third, for each day $t$, we calculate the fraction of hourly-bins spent at each location during nighttime (midnight to 6 a.m.
+for home detection), and typical business hours (9 a.m.
+to 4 p.m.
+on weekdays, for work detection).
+This choice is motivated by the observation that individuals are most likely at home between midnight and 6 a.m.
+($\sim 75\%$ probability) and at work between 9 a.m.
+and 4 p.m.
+on workdays ($\sim 50\%$), see Section 4.
 To compute such a fraction, we use as a denominator the number of hourly bins with records (Figure 2.b).
@@
 Fourth, for each day $t$, we consider a window of size $\Delta_{T}+1$, covering the range included in $t\pm\Delta_{T}/2$ days (see Figure2.c).
-This window acts as a memory component for the algorithm to better detect behavior in the period before and after a specific day.
+A configurable parameter allows restricting the window to past data only, corresponding to the interval $t-\Delta_{T}$ days (see Supplementary materials Sec.
+S-8).
+This window acts as a memory component for the algorithm to better detect behaviour in the period before and after a specific day.
 In each window, we average the fraction of hours a location was visited (see Figure 2c).
@@
 Among these, we retain only those that meet minimum requirements with respect to the proportion of visits within and across days.
-These requirements are controlled by parameters $dn_{H,W}$, $hf_{H,W}$, $df_{H,W}$ (see Table 2).
+These requirements are controlled by parameters $C_{days}$, $f_{hours}$, $f_{days}$ (see Table 1).
 If no location meets the requirements, no home (or work) is assigned for that specific window.
@@
 The top location from these two lists is then assigned as home (or work).
-By prioritizing the fraction of days over the fraction of hours for work detection, we are emphasizing work as a location defined by regular visits throughout the workweek (see Sec.SI-6).
-In summary, the algorithm allows the control of two aspects.
-It enables the user to specify the criteria for determining when there is enough data to attempt home and work detection.
-Also, it enables to control of when a location can be considered as a candidate home or work location, all while accounting for the fact that users can have different levels of missing data.
+By prioritising the fraction of days over the fraction of hours for work detection, we are emphasising work as a location defined by regular visits throughout the workweek (see Supplementary materials Sec.
+S-8).
+In summary, the algorithm offers control over two key aspects.
+It allows users to define the threshold for determining when there is enough data to attempt home or work detection.
+It also enables users to set the criteria for determining whether a location qualifies as a candidate home or work site, while accounting for individual differences in data completeness.
 Figure 2: Schematics of the HoWDe algorithm workflow for home/work location detection.
@@
 The pie charts display the allocation of visits across locations during the night hours (orange) and work hours (blue) on each day.
-Note that on Saturday, 6/01, the night hours data was discarded since there were not enough hours with available data (exemplifying the filter tuned by the parameter $dhn$).
+Note that on Saturday, 6/01, the night hours data was discarded since there were not enough hours with available data (exemplifying the temporal coverage filter tuned by the parameter $C_{hours}$).
 In this minimal example, no location appears during both night and work hours; however, this may happen in real data.
 c.
-For each day, $t$ we consider a sliding window of $3$ consecutive days centered on $t$.
+For each day, $t$ we illustrate a sliding window of 33 consecutive days centered on $t$.
 For work detection, the sliding window excludes weekends.
 d.
-Aggregation step for the window centered on Friday, 5th.
+Aggregation step for the window centred on Friday, 5th.
 For each location, we compute the average fraction of time it was visited during night hours (orange) and during work hours (blue).
 Additionally, we compute the fraction of days the location is visited at least once during work hours.
-Locations are then sorted in descending order based on these fractions, and the top ones are selected as the estimated Home (L1) and Work (L2) locations on Friday.
-2.3 Validation against ground truth data
-To assess how well the procedure described above can capture individuals’ home and work locations, we measure i) the accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Methods for more details).
-We run the validation using the ground-truth datasets D1 and D2, described in Methods.
+Locations are then sorted in descending order by on these fractions, and the top ones are selected as the estimated Home (L1) and Work (L2) locations for Friday.
+5.2 HoWDe configurable parameters
+To adapt to diverse datasets and research goals, HoWDe includes a set of tunable parameters that govern the algorithm’s temporal coverage filtering and the temporal consistency required to assign home and work locations.
+First, we define the size of the memory window to be configured separately for home and work $\Delta\mathbf{T_{H}}$, $\Delta\mathbf{T_{W}}$.
+Second, the coverage parameters specify the minimum temporal coverage required at the daily ($C_{hours}$) and the window level ($C_{days,H}$, $C_{days,W}$):
+  • $\mathbf{C_{hours}}$: This parameter specifies the minimum fraction of night- or business-hour bins that must contain data for a day to be considered in the detection process.
+By default, HoWDe retains only days with at least 40% of the hourly bins containing stop information, ensuring a minimum level of temporal coverage.
+The threshold can be adjusted depending on the data quality or the application context.
+  • $\mathbf{C_{days,H}}$ and $\mathbf{C_{days,W}}$: To account for variability in data completeness across datasets, HoWDe requires a minimum fraction of days with data within each sliding window.
+These parameters independently control home and work detection.
+If the fraction of days with data in the window falls below the threshold, the algorithm marks the day as having an “undetected” home or work location.
+Third, the location-selection parameters determine when a place qualifies as a potential home or work location, based on the minimum fraction of hours within a day ($f_{hours,H}$, $f_{hours,W}$) and the minimum fraction of days within a window ($f_{days,W}$):
+  • $\mathbf{f_{hours,H}}$ and $\mathbf{f_{hours,W}}$: These parameters set the minimum average fraction of hourly bins (night or business hours, respectively) during which a candidate location must be visited to qualify as home or work.
+If the location’s average fraction of hourly-bin visits over the window falls below this threshold, it is excluded as a potential home or work location.
+  • $\mathbf{f_{days,W}}$: Work locations often follow different visit regularities than home locations, reflecting local labour norms (e.g., four-day work weeks).
+To capture these patterns, HoWDe requires that a work location be visited on at least a minimum fraction of days with data within the window.
+This parameter does not apply to home detection, which typically assumes near-daily presence when data are available.
+Table 1: Summary description of the HoWDe configurable parameters.
+The table reports the parameter names, whether they are used in home detection, work detection, or both, and the description of the parameter’s purpose.
+When looking at the hourly bins with data (hours with data), that is the case of $C_{hours}$, $f_{hours,H}$, and $f_{hours,W}$, fractions are computed with respect to night/business hourly bins only.
+Note that $C_{hours}$, $C_{days,H}$, and $C_{days,W}$ act as temporal coverage filters to ensure a given day has enough data to estimate home or work location.
+Instead, $f_{hours,H}$, $f_{hours,W}$, and $f_{days,W}$ focus on location-specific visit requirements, allowing each location above the threshold to be a candidate home/work location.
+Specifically, fractions of visits to home/work location candidates are computed relative to the total hours in which an individual has recorded data during the specified period.
+Name Usage Description
+$\Delta T_{H}$ Home estimation Size of the memory window for home estimation
+$\Delta T_{W}$ Work estimation Size of the memory window for work estimation
+$C_{hours}$ Home/Work temporal coverage filter Min. fraction of night/business hourly-bins with data in a day
+$C_{days,H}$ Home temporal coverage filter Min. fraction of days with data in a window
+$C_{days,W}$ Work temporal coverage filter Min. fraction of days with data in a window
+$f_{hours,H}$ Home location selection Min. fraction of night hourly-bins a location should be visited
+$f_{hours,W}$ Work location selection Min. fraction of business hourly-bins a location should be visited
+$f_{days,W}$ Work location selection Min. fraction of days a location should be visited within the window
+For details on parameter implementation, we refer to the Supplementary material Section S-8, where we provide the algorithm’s pseudocode.
+6 Results
+Here, we evaluate the performance of HoWDe against ground truth data, explore the effect of parameter choices, and assess its robustness across demographic groups.
+Then, we demonstrate how HoWDe can be applied to estimate employment rates and commuting patterns.
+6.1 Algorithm performance: accuracy, retention, and robustness
+To assess how well HoWDe can capture individuals’ home and work locations, we measure i) the detected accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Sec.
+3.3 for more details).
+We run the validation using the ground truth datasets D1 and D2, described in Methods.
 In dataset D2, annotators reviewed complete stop sequences; therefore, the temporal dimension (changes of home and work locations) can not be considered.
-For this reason, for D2, we consider a single window covering the entire data period (see Sec.SI-7 for parameter details).
+For this reason, for D2, we consider a single window covering the entire data period (see Supplementary material Sec.
+S-9 for parameter details).
 6.1.1 Algorithm performance across parameter choices
-We find that sliding window length $\Delta_{T}$ and minimum home-time fraction $hf_{H}$ have the greatest impact on the accuracy of HoWDe’s home detection, while other parameters have a smaller effect (see Sec.SI-8).
-Both the accuracy of the home location detection and the fraction of non-detected homes increase as the minimum home-time fraction threshold $hf_{H}$ increases (Figure 3a,b).
-In dataset D1, the fraction of non-detected homes increases with the window size, but accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks).
+We evaluate how the choice of configuration parameters affects HoWDe’s performance.
+Among all parameters, the sliding-window length $\Delta_{T}$ and the minimum home-time fraction $f_{hours}$ have the strongest influence on home-detection accuracy, while other parameters exert more moderate effects.
+Temporal coverage filters, in particular, primarily determine the amount of usable data and should therefore be tuned according to the quality and completeness of the dataset at hand.
+A detailed analysis of these additional parameter configurations, including their dependence on the datasets’ sparsity, is provided in Supplementary Material S-10.
+Both the detected accuracy of the home location estimation and the fraction of non-detected homes increase as the minimum home-time fraction threshold $f_{hours,H}$ increases (Figure 3a,b).
+In dataset D1, the fraction of non-detected homes increases with the window size, but the detected accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks).
 This suggests that a window of $4$ weeks offers a good balance between high accuracy and user retention for home detection.
 Figure 3: Home and Work validation across configurations a.
-Accuracy of the home location detection for datasets D1 (first column) and D2 (second column).
-Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night.
+Detected accuracy of the home location detection for datasets D1 (first column) and D2 (second column).
+Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night ($f_{hours,H}$).
 For dataset D1, we explore how the results change for each sliding window increase (x-axis), while for dataset D2, we focus on a single sliding window (comprising the entire data period).
 b.
-Fraction of not detected of the home location detection for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend.
+Fraction of not detected home locations for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend.
 c.
-Accuracy of the work location detection for datasets D1 (first column) and D2 (second column).
-Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours, and the minimum fraction of days an individual should visit the potential work across weekdays.
-For dataset D1, we explore how the results change for each sliding window increase, while for dataset D2, we keep the same sliding window size (comprising the entire period).
+Detected accuracy of the work location detection for datasets D1 (first column) and D2 (second column).
+Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours ($f_{hours,W}$), and the minimum fraction of days an individual should visit the potential work across weekdays ($f_{days,W}$).
+For dataset D1, we explore how results change with increasing sliding window sizes, while for dataset D2, we keep the same sliding window size (covering the entire period).
 d.
-Fraction of not detected of the work location detection for dataset D1 (first column) and dataset D2 (second column).
+Fraction of not detected work locations for dataset D1 (first column) and dataset D2 (second column).
 The legend is shared with panel c.
-Table 1 reports the results by focusing on two configuration setups: the minimum configuration, which maximizes accuracy while minimizing the number of non-detected users, and the maximum configuration, which prioritizes maximum accuracy at the cost of higher user non-detection rates.
-We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [28, 2] and TimeGeo [29] (see Methods for more details on these algorithms).
-The maximum configuration is reached for $\Delta_{T}=28$ days window and $90\%$ minimum home-time fraction (see Sec.SI-7 for parameter details).
-With this choice of parameters, the algorithm achieves an accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 1).
-The minimum configuration is obtained for parameters $\Delta_{T}=28$ days window and $50\%$ home frequency (see Sec.SI-7 for parameter details).
-Here, the algorithm achieves $92.94\%\pm 0.49\%$ accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
-All in all, we find that HoWDe achieves higher accuracy than previous heuristic methods [29, 2], partly because it excludes users with low-quality data (see Table 1).
-Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T}=42$days ($t\pm 3$ weeks), after which accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
-In the work location detection, two key parameters are the minimum fraction of hourly bins ${hf_{W}}$ and days ${df_{W}}$ required for a location to be selected as work.
-Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance accuracy in D1 but increase non-detection rates.
-For a larger value of ${hf_{W}}=60\%$, increasing ${df_{W}}$ enables significant accuracy gains.
-In contrast, for ${hf_{W}}=40\%$, changes in ${df_{W}}$ have a weaker effect.
-We find similar results for D2, when changing the minimum fraction of days ${hf_{W}}$ and fraction of hours ${hf_{W}}$.
-HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Sec.SI-7 for parameter details), outperforming baseline methods (see Table 1).
-Similarly, for D2, HoWDe achieves high accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 1).
-This difference in performances may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
-Accuracy HoWDe min HoWDe max Atlas TimeGeo
-Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $81.96\pm 0.55$ $79.28\pm 0.78$
-D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.77\pm 1.21$ $70.38\pm 1.12$
-Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $60.32\pm 0.80$ $39.51\pm 0.92$
-D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $69.71\pm 2.17$ $57.04\pm 2.40$
+Table 2 reports the results by focusing on two configuration setups: the minimum configuration, which maximises detected accuracy while minimising the number of non-detected users, and the maximum configuration, which prioritises maximum detected accuracy at the cost of higher user non-detection rates.
+We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [32, 3] and TimeGeo [33] (see Secs.
+2 and 3.4 for algorithm and performance comparison details).
+The maximum configuration is reached for $\Delta_{T_{H}}=28$ days window and $90\%$ minimum home-time fraction (see Supplementary material Sec.
+S-9 for parameter details).
+With this choice of parameters, the algorithm achieves a detected accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 2).
+The minimum configuration is obtained for parameters $\Delta_{T_{H}}=28$ days window and $50\%$ home frequency (see Supplementary material Sec.
+S-9 for parameter details).
+Here, the algorithm achieves $92.94\%\pm 0.49\%$ detected accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
+All in all, we find that HoWDe achieves higher detected accuracy than previous heuristic methods [33, 3], partly because it excludes users with low-quality data (see Table 2).
+Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T_{W}}=42$ days ($t\pm 3$ weeks), after which detected accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
+In the work location detection, two key parameters are the minimum fraction of hourly bins $f_{hours,W}$ and days $f_{days,W}$} required for a location to be selected as work.
+Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance detected accuracy in D1 but increase non-detection rates.
+For a larger value of ${f_{hours,W}}=60\%$, increasing $f_{days,W}$} enables significant accuracy gains.
+In contrast, for ${f_{hours,W}}=40\%$, changes in $f_{days,W}$} have a weaker effect.
+We find similar results for D2, when changing the minimum fraction of days $f_{hours,W}$} and fraction of hours $f_{hours,W}$}.
+HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Supplementary material Sec.
+S-7 for parameter details), outperforming baseline methods (see Table 2).
+Similarly, for D2, HoWDe achieves high detected accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 2).
+This difference in performance may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
+While Atlas does not explicitly estimate work locations, our adaptation (see Sec.
+2) outperforms TimeGeo.
+This highlights that methods based solely on distance and visit counts, such as TimeGeo, perform substantially worse than those incorporating visit duration, including our Atlas adaptation and HoWDe.
+By accounting for missing data within individuals, HoWDe achieves the highest overall detected accuracy.
+Detected accuracy HoWDe min HoWDe max Atlas TimeGeo
+Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $82.92\pm 0.66$ $81.44\pm 0.70$
+D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.62\pm 1.06$ $70.54\pm 1.15$
+Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $62.20\pm 0.92$ $37.51\pm 0.92$
+D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $71.74\pm 2.17$ $56.93\pm 2.35$
 Fraction not detected HoWDe min HoWDe max Atlas TimeGeo
-Home D1 $4.29\pm 0.38$ $38.75\pm 0.93$ $4.60\pm 0.31$ $\mathbf{0.09\pm 0.06}$
+Home D1 $4.29\pm 0.38$ $38.75\pm 0.93$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
 D2 $10.44\pm 1.02$ $39.94\pm 1.43$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
-Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $3.45\pm 0.31$ $\mathbf{1.86\pm 0.23}$
-D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.67\pm 0.41$
-Table 1: Evaluation of home and work detection across methods and datasets.
-2.4 Robustness and biases across configurations
+Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $\mathbf{0.0\pm 0.0}$ $0.80\pm 0.16$
+D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.68\pm 0.36$
+Table 2: Performance comparison of home and work detection methods.
+Detected accuracy (top) and fraction not detected (bottom) across datasets D1 and D2.
+Bold highlights indicate the best method per metric—highest accuracy or lowest missing rate.
+While Atlas does not explicitly detect work locations, we adapt its home location algorithm for this purpose (see Section 2).
+6.1.2 Algorithm robustness across demographic groups
 HoWDe’s accuracy comes at the cost of user retention.
 The exclusion of users may introduce biases toward specific demographic groups.
-To assess these biases, we measure accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanization levels, and country GDP.
-Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Sec.SI-7).
-The largest accuracy gaps appear in suburban areas and middle-income countries: home detection drops by $3.39$ points in suburban areas (compared to urban areas) and work detection by $9.13$ in middle-income countries (compared to high-income countries).
-Despite accuracy deviations, home detection accuracy remains within $90\%$ $99\%$ and work detection within $77\%$ $89\%$ across all lower-accuracy subgroups (see Sec.
-SI-9), aligning with the observed range for D1 and D2.
-Nevertheless, even minor detection errors can propagate, compromising the overall reliability of the results.
+To assess these biases, we measure detected accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanisation levels (estimated using the 2022 GHS settlement model grid at 1 km resolution [62]), and country GDP.
+Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Figure 4), and explored only for subgroups comprising at least 200 users.
+The largest detected accuracy gaps appear in suburban areas and middle-income countries.
+In suburban settings, home detection accuracy drops by $3.39$ points compared to urban areas under the minimum configuration.
+However, user retention remains stable, suggesting that the decline in detected accuracy is not due to a disproportionate loss of suburban users.
+In middle-income countries, home detection accuracy falls only under the maximum configuration (3.31-3.31 points), while work detection exhibits larger drops of $9.13$ and $8.06$ points under the maximum and minimum configurations, respectively.
+These declines occur without changes in user retention, suggesting that the reduced detected accuracy reflects differences in work visitation patterns rather than sample loss.
+Across other demographic groups, we find significant differences in home detection.
+Home detected accuracy drops by $1.87$ points for males compared to females under the minimum configuration.
+For individuals aged 20–30, detected accuracy decreases by $2.40$ and $3.60$ points under the minimum and maximum configurations, respectively, relative to the 40–60 age group.
+Despite accuracy deviations, home detection accuracy remains within $90\%-99\%$ and work detection within $77\%-89\%$ across all lower-accuracy subgroups (see Figure 4), aligning with the observed range for D1 and D2.
+These findings emphasize that, while detection accuracy remains consistent across demographic groups and aligns with overall results, it is essential to validate sample representativeness when applying home and work detection algorithms.
+Figure 4: Home and work detection accuracy and fraction of not detected locations across demographics for dataset D1.
+a.
+Detected accuracy of home location detection across demographic groups.
+Dark orange illustrates the results for the maximum configuration, while light orange illustrates the minimum.
+Error bars indicate bootstrapped error estimates (applies to all panels).
+b.
+Fraction of not detected home locations (configurations color coding is shared with panel a).
+c.
+Detected accuracy of work location detection across demographic groups.
+Dark blue illustrates the results for the maximum configuration, while light blue illustrates the minimum.
+d.
+Fraction of not detected work locations (configurations color coding is shared with panel c).
+6.2 Impact of parameter sensitivity on real-world applications
+Small detection errors can propagate across applications, potentially undermining the reliability of downstream results.
 To address this potential issue, we examine how tuning the detection algorithm parameters impacts results in two downstream applications: (i) employment rate estimates in Europe and (ii) commuting pattern characteristics.
@@
 However, while both exhibit moderate to high Pearson correlation, the maximum configuration has a stronger correlation with reported employment rates but estimates fewer regions due to stricter data coverage thresholds.
-It also has a higher overall relative error (Figure 4.b, and Figure SI-12, except in Germany, where the trend reverses.
+It also has a higher overall relative error (Figure 5.b, and Supplementary material Figure S-18, except in Germany, where the trend reverses.
 This suggests that while the maximum configuration enhances individual accuracy, the minimum configuration may be more suitable for population estimates.
@@
 This case study examines whether commuting distance, measured as the great-circle distance between detected home and work locations, varies with population density and whether parameter configurations can influence results.
-We estimate urbanization levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [40].
+We estimate urbanisation levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [62].
 Our analysis reveals consistent findings across configurations: commuting distances are longer in rural areas compared to urban areas, with an absolute difference of $3.61\pm 0.04$ km for the “minimum” and $3.85\pm 0.06$ km for the “maximum” configurations (see Figure 5.c).
@@
 Figure 5: Applications for home and work detection a.
-Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest accuracy (right).
+Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest detected accuracy (right).
 b.
@@
 7 Discussion
-In this work, we contributed to the development of accurate and robust methods for home and work detection.
-Using two datasets (D1, D2) covering $\sim 5,100$ individuals from over $80$ countries, we characterized home/work visit patterns.
-We found that individuals are most likely at home between midnight and 6 AM ($\sim 75\%$ probability) and at work between 9 AM and 4 PM on workdays ($\sim 50\%$).
+Home and work location detection from smartphone data is a widely used task in mobility and transport research.
+Despite its central role in many studies, most existing approaches have not been systematically validated against ground truth and do not explicitly account for the heterogeneous and irregular patterns of missing data that characterise real-world smartphone datasets.
+In this work, we proposed HoWDe, an algorithm for home and work location detection from stop-location data.
+HoWDe integrates key features of state-of-the-art heuristic approaches and improves robustness against missing data by allowing for heterogeneous data gaps within and across users.
+HoWDe builds on insights derived from empirical home-work visit data.
+Using two datasets capturing stop-locations and ground truth home and work locations (D1, D2) for $5,099$ individuals from $68$ countries, we characterised home/work visit patterns.
+We found that individuals are most likely at home between midnight and 6 a.m.
+($\sim 75\%$ probability) and at work between 9 a.m.
+and 4 p.m.
+on workdays ($\sim 50\%$).
 However, visit patterns and missing data vary widely across users and over time, and even within individual users’ routines.
-Building on these insights, we provided an open-source implementation of HoWDe, a home/work detection algorithm that integrates existing approaches while improving robustness against missing data.
-We validated HoWDe against ground truth datasets, achieving up to $\sim 97\%$ accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ accuracy in D1 and $\sim 82\%$ accuracy in D2 for work detection.
-Our analysis highlights the trade-off between accuracy and retaining users with low-quality data.
-We identified two key parameters affecting performance: the window length considered and the minimum visit proportion required for classification.
+We validated HoWDe against the ground truth datasets, achieving up to $\sim 97\%$ detected accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ detected accuracy in D1 and $\sim 82\%$ detected accuracy in D2 for work detection.
+We found that Atlas and Timegeo achieve lower overall detected accuracy than HoWDe.
+However, they attain high home-detection performance ($81.44\%-82.92\%$ for D1 and $70.10\%-70.62\%$ for D2), indicating that both are adequate for this task.
+In contrast, for work detection, relying solely on distance and visit counts—as in Timegeo—yields much lower detection accuracy ($37.51\%$ for D1 and $56.89\%$ for D2), whereas incorporating duration, as in our adaptation of Atlas and in HoWDe, results in markedly higher accuracy.
+Our analysis highlights the trade-off between achieving accuracy and retaining users with low-quality data.
+We identified two key parameters affecting performance: the window length considered ($\Delta_{T}$) and the minimum visit proportion required for classification ($f_{hours,H,W}$, $f_{days,W}$).
 A window of at least $28$ days ensures a reliable home estimation, while work windows longer than $42$ days reduce user retention without significant accuracy gains.
-We also revealed how the algorithm’s accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection).
+We also revealed how the algorithm’s detected accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection).
 Although these accuracy gaps persist across configurations, two case studies show that prioritizing accuracy over user retention is preferable for individual-level tasks, whereas maximizing retention may be more valuable for population-level analyses.
@@
 Despite its strengths, HoWDe has key limitations.
-While it allows customization of detection windows and thresholds, more advanced methods may be required to fully capture mobility patterns that are collected by diverse GPS-equipped devices [43].
-Its performance may also vary with lower temporal resolution or geographically limited datasets, particularly in underrepresented low-income regions with sparser mobile data.
-Future work should assess its applicability in rural and less densely populated areas [44].
-Detecting home and work locations raises important ethical concerns, particularly regarding privacy [45, 46], as the geographical coordinates of these locations are highly identifying and can enable re-identification of individuals in mobility datasets.
-Currently, many data holders apply in-house heuristics to detect home and work, but these methods are often opaque, introducing methodological uncertainty.
-By providing a validated and open-source detection framework, our approach enables data holders to perform this sensitive step internally using a standardized and transparent pipeline.
-We hope this will facilitate the sharing of privacy-preserving datasets—for example, through the release of obfuscated or aggregated representations—while ensuring methodological consistency and scientific utility for downstream research.
+While providing a robust framework for detecting anchor locations, it does not infer their semantic purpose—that is, it cannot distinguish between work, study, or other activities with similar temporal patterns.
+Future work could address this by integrating contextual data, such as point-of-interest (POI) information or survey-based semantic labels.
+In addition, HoWDe assumes a single dominant lifestyle pattern across users and therefore cannot, within a single run, distinguish between regular schedules and rotating night-shift routines.
+While this assumption aligns with the empirical regularities in our ground truth data, incorporating an initial lifestyle-detection step represents a promising direction for future work.
+Finally, the current implementation identifies a single home and work location per day and does not explore multi-home or multi-work scenarios.
+Leveraging the ranked lists of potential home and work locations (stored internally in HoWDe) could support such analyses, but this lies beyond the evaluation possibilities afforded by our ground truth data.
+More advanced methods may also be needed to handle heterogeneous GPS data and ensure its applicability across contexts, in particular for rural and less densely populated areas [41, 65].
+Working with individual-level location data presents well-known challenges for data sharing and reproducibility due to the high risk of re-identification [66, 67].
+Under the EU General Data Protection Regulation, re-identification risk arises when: (i) the data is not k-anonymous—that is, individual trajectories are unique—and (ii) the data can be linked, through means that are reasonably likely to be used, to an external dataset containing direct personal identifiers (GDPR Recital 26).
+The first condition is almost always met in human mobility datasets: even after coarse-graining or obfuscating data points, trajectories remain highly idiosyncratic [66], making k-anonymity infeasible without rendering the data analytically useless.
+The second condition can be mitigated through privacy-preserving techniques—such as perturbing spatial coordinates, obfuscating timestamps, or omitting sensitive fields—to increase the difficulty of linking mobility traces to known individuals.
+These transformations can in some cases reduce re-identification risk to a negligible level, but often at the cost of limiting the usefulness of the data for certain analytical purposes.
+This creates a fundamental trade-off between privacy protection and scientific utility, which is why access to individual-level location data is typically restricted to secure environments and regulated by data-use agreements.
+Because of these constraints, many data holders perform sensitive inference steps—such as detecting home and work locations—internally.
+However, the methods used for this purpose are often proprietary or undocumented, introducing uncertainties and reducing reproducibility.
+Our contribution directly addresses this gap by providing a validated, open-source pipeline for detecting home and work locations that can be executed in-house by data holders prior to data sharing.
+This supports methodological transparency and consistency, while enabling the release of privacy-preserving data products—such as aggregated visitation patterns or synthetic traces.
 In summary, by advancing methods for home and work detection, this study contributes to a more robust and scalable foundation for mobility research.
 By providing an open-source, validated solution, we hope to facilitate more reliable and ethical research, contributing to a deeper understanding of human mobility and its implications for society.
-4 Material and Methods
-4.1 Dataset 1: DTU stop data and self-reported annotations
-The first dataset re-uses data collected during 2017-19 already used in previous publications [3], where stop locations were detected using the scalable stop-location detection algorithm InfoStop [47].
-The stop location data captures the mobility patterns of 96,065 users between 25 and 64 years old for 81 countries across continents.
-The dataset also contains self-recorded home and work labels for $2,309$ labelled home and $2,479$ work locations, collected across $4,599$ mobile phone and smart band users.
-Note that the labelled dataset is biased towards urban areas and higher- to middle-income countries.
-The spatial coordinates of locations are no longer available to preserve individuals’ privacy.
-All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).
-4.2 Dataset 2: Veraset Movements data and annotation
-Anonymous raw GPS points are provided at a global level by the Veraset data company in the framework of the World Bank “Monitoring COVID-19 policy response through human mobility data” project.
-Veraset Movements data consists of raw GPS point information from devices all around the world.
-In contrast to the majority of GPS data providers, Verasets Movements are sourced from thousands of Software Development Kits (SDK) to reduce sample biases.
-It features coverage of about $5\%$ of the global population [48].
-In our work, unique-device GPS points are processed and analyzed as a proxy of individuals’ mobility and patterns of visits.
-The data at our disposal spans over a period of $12$ months, starting from Jan 01, 2020, until Dec 31, 2020.
-In this work, multiple countries were analyzed to assess the impact of the pandemic and the policy responses across a vast spectrum of potentially different behaviors.
-Dataset D2, in particular, consists of the GPS trajectories of $100$ distinct devices for each of the $5$ selected countries: Brazil, Colombia, Indonesia, Mexico, and the Philippines.
-For each country independently, we require $33\%$ of the users to be living in high-wealth neighbourhoods, $33\%$ of the users to be living in medium-wealth neighbourhoods, and the remaining part in low-wealth neighbourhoods.
-Wealth classification and users’ place of residence are drawn leveraging previous work on GPS mobility in middle-income countries [15].
-In particular, each user is also required to have a minimum level of recorded activity, consisting of at least a stop record in $20\%$ in the days before the COVID-19 pandemic declaration and $20\%$ of days with at least a stop record over the entire data period.
-This second dataset involves, for each user’s stop location sequences, the expert annotation through detailed inspection of location visits, satellite imagery, and local amenities, focusing on five middle-income countries (Brazil, Colombia, Indonesia, Mexico, and the Philippines) for 500 individuals, from which 287 have at least one work location annotated (see Sec.
-SI-1 detailed description of the annotation procedure).
-Given that annotators only labelled the complete longitudinal stop sequences, an explicit consideration of changes in home and work for each individual is not possible.
-4.3 Home and Work locations detection algorithms: a primer
-Smartphones use an in-built GPS sensor to estimate the device’s latitudes and longitude coordinates over time, at a frequency that varies depending on device usage.
-The device position is then collected by various location-based apps, such as navigation or social media apps.
-These data may then be used by location intelligence companies or researchers to study human mobility and its regularities, both at the individual and at the population level.
-Raw location data utilized for mobility analysis consists of timestamped sequences of (latitude, longitude) records associated with a unique user ID.
-The first pre-processing step involved in the analysis is the so-called “stop-location detection” [47, 49].
-Stop location detection consists of (i) identifying records that are captured when individuals stay within a small area for a sufficient amount of time (e.g.
-when they are at home, at the office, or in a restaurant); (ii) applying spatial clustering to all these “static records”, to identify a set of unique locations, i.e.
-those places where people spend time.
-After detecting the individuals’ stops-locations, individual trajectories can be summarized as sequences of stays.
-Each stay is associated with a start time (when the stay started), a duration (how long the stay lasted), and a location ID (where the stay happened).
-A second critical processing step often involved in mobility studies is to identify key locations, such as individuals’ home and work locations, starting from the sequences of stays.
-To achieve this, researchers and practitioners often define heuristics based on the assumption that visits to home and work locations follow specific daily and weekly patterns.
-While a standard practice does not exist, in this work, we focus on two distinct and widely used state-of-the-art methods that exploit travel circadian rhythms and pattern regularities.
-These methods were designed and tested over the past decade with different approaches, in a collective scientific effort, separately leveraging diverse sets of mobile-phone-based data (e.g.
-Call Detail Records (CDRs), eXtended Detail Records, (XDRs), or GPS records) [21, 50, 51, 35] as well as activity-based travel surveys [36].
-The main rationale is to rank locations based on the day of the week (weekdays vs.
-weekends) and the time of day (daytime vs.
-nighttime) when visits occur.
-4.3.1 The Atlas approach
-The method used in [2, 6, 52, 28] (hereinafter “Atlas” method) defines a home location as the most probable location visited on weekdays between 10 p.m.
-and 6 a.m., i.e.
-the stop location where a device was found spending most of the weekday nighttime.
-While the Atlas methods mostly focuses on home locations, its definition can be easily generalized to the case of work locations.
-Following a similar rationale, we define an Atlas work location as the most probable stop location visited during weekdays daytime.
-4.3.2 The Time-Geo approach
-In contrast with the definition introduced by the Atlas method, more complex heuristics are adopted in Jiang et al.
-[29] (hereinafter referred to as the “Time-geo” method), in particular for the case of identifying work locations.
-Time-geo ranks stop locations based on the frequency of visits during weekday nights (between 7 p.m.
-and 8 a.m.) and weekends and picks the most visited location as an individual’s home.
-The same logic is applied when detecting the work location, but now focusing on weekday work time boundaries (between 8 a.m.
-and 7 p.m.) and including additional requirements: i) the candidate’s work location is visited at least three times across the study period, and ii) the distance between the work candidate and an individual’s home location is greater than $500$ meters.
-Note that both methods rely on the definition of stop location.
-Stop location definition might vary depending on the type of data analyzed and the specific characteristics of the data points, e.g.
-whether you use Call Detail Records (CDR), eXtended Detail Records (XDR), or GPS points.
-However, the principled choices these algorithms are founded on can naturally be extended to any of these different cases, making them easily generalizable.
-In this work, stop locations are defined as spatiotemporally close sets of GPS points.
-All comparisons between home-work detection methods are always drawn from a set of stop locations constructed following identical procedures.
-For more information about stop location processing, we refer to the Supplementary Information (SI) Sec.SI-2 and the existing literature [47, 14].
-4.4 Home and work visits behavioral clusters
-Section 2.1 focused on the temporal attributes of users’ daily mobility patterns.
-To better highlight groups of users with similar behavioral features, we reported the results in Figure  1.c-f, sorting daily patterns based on the visit hour of the first home visit.
-behavioral groups were generated using the K-Modes clustering algorithm, a categorical adaptation of the traditional K-Means algorithm [53].
-Here, we use daily sequences of 24 hourly bins with three possible stop categories.
-For example, in the case of home patterns, each sequence consists of either a home (category 1), a visit to “a different location”, e.g., work or recreational locations (category 2), or an hour with missing data (category 3).
-A similar approach is used for work patterns, where visits to work location are considered as “category 1”, while visits to home location are considered as “a different location” (category 2).
-The K-Modes algorithm requires as a parameter, the number of desired clusters to compute.
-In Figure 1.c,e we reported 6 clusters.
-In Figure 1.d,f we reported 3 and 4 clusters, respectively.
-Note that, while the number of clusters is the result of the selection of cluster number following the elbow method for each of the four plots, our purpose is to simply highlight the main different visit patterns.
-Following this idea, in Sec.SI-3, where we present results for the smaller D2 dataset, we manually selected three as the number of clusters used to report the hourly day-visit patterns.
-4.5 Behavioral profile entropy
-We quantify the degree to which users transition between behavioral profiles measuring the normalized Shannon entropy [38], treating each behavioral cluster as a distinct state, following:
-$H_{\text{norm}}=-\frac{1}{\log{k}}\sum_{i=1}^{k}p_{i}\log{p_{i}},$ (1)
-where $k$ is the total number of possible clusters, and $p_{i}$ is the probability of occurrence of each state, calculated as: $p_{i}=\frac{n_{i}}{N}$ where $n_{i}$ are the total days classified in cluster $i$ and $N$ the total number of days for a given user.
-Normalizing by the maximum entropy $\frac{1}{\log{k}}$, ensures that entropy remains comparable across cases (weekend vs.
-weekday and home vs.
-work), staying within the same range from 0 (when there is only one profile) to 1 (when there are multiple profiles, and all are equally important).
-4.6 HoWDe parameters details
-$\mathbf{dhn}$: First, to account for a minimum level of information within a day, we focus only on days with hour sequences consisting of at least three stops, i.e.
-stops visited in at least three hours of the day (i.e.
-days with at least three hours of data).
-The HoWDe algorithm, however, can be tuned to handle different thresholds based on the specific needs or constraints in place.
-$\mathbf{dn_{H,W}}$: Second, to account for the potentially varying data quality across different datasets, we introduce two independently configurable pairs of parameters for home and work locations.
-The first pair of parameters works as a maximum fraction of days without data per sliding window, marking a day with an “undetected home/work location” if the fraction of days without data within the sliding window exceeds the threshold.
-$\mathbf{hf_{H,W}}$: A third pair of parameters threshold the minimum fraction of hours per night or day period for a location to be home or work respectively.
-If the average fraction of visits across all days in the sliding window is below the threshold, the location cannot be considered a potential home/work.
-$\mathbf{df_{H,W}}$: Fourth, while home locations are expected to be visited almost every day (provided that enough data are available), work locations are likely to follow a different pattern, as those locations will be subject to local work habits, e.g.
-four days working week.
-To account for regularities across days, we introduce an additional requirement for work locations: a work location should be visited for a minimum fraction of days with data.
-Table 2: Summary description of the HoWDe configurable parameters.
-The table reports the parameter names, whether they are used in home detection, work detection, or both, and the description of the parameter purpose.
-When looking at the hourly bins with data (hours with data), that is the case of $dhn$, $hf_{H}$, and $hf_{W}$, counts and fractions are computed with respect to night/business hourly bins only.
-Note that, $dhn$, $dn_{H}$, and $dn_{W}$ act as minimum activity thresholds for a given day to have an estimated home or work location estimated.
-$hf_{H}$, $hf_{W}$, and $df_{W}$ serve a similar purpose but for location-specific visit requirements allowing each location above the threshold to be a candidate home/work location.
-Specifically, fractions of visits to home/work location candidates are computed relative to the total hours in which an individual has recorded data during the specified period.
-Name Usage Description
-$\mathbf{\Delta_{T}}$ Home/Work data quality Size of the memory window
-$\mathbf{dhn}$ Home/Work data quality Min. number of night/business hourly-bins with data in a day
-$\mathbf{dn_{H}}$ Home data quality Max. fraction of days without data in a window
-$\mathbf{dn_{W}}$ Work data quality Max. fraction of days without data in a window
-$\mathbf{hf_{H}}$ Home estimation Min. fraction of night hourly-bins (avg. over days in the window)
-$\mathbf{hf_{W}}$ Work estimation Min. fraction of business hourly-bins (avg. over days in the window)
-$\mathbf{df_{W}}$ Work estimation Min. fraction of days within the window
-For details on parameter implementation, we refer to the Sec. SI-6, where we provide the algorithm’s pseudocode.
-4.7 Accuracy and “fraction of not detected” definition
-In this work, we report the performance of the algorithm in terms of two main quantities: “accuracy” and “fraction of not detected”.
-Here, we provide the definitions of these two concepts.
-The definitions below are explained in the case of the home detection, and they apply in the same way for the work detection.
-The fraction of ‘not detected’ indicates cases where a home location was labelled, but the algorithm couldn’t assign a home due to the thresholds on data quality or minimum fraction of visits.
-In the case of D1, the dataset consists of a set of user-weeks, corresponding to users that self-reported living in a specific home (the “ground truth home”) in a given week.
-For some of these user-weeks, our algorithm identifies a “detected home”.
-We evaluate the performance of the algorithm by comparing pairs of “ground truth home” and “detected home” across user-weeks.
-The “fraction of not detected” is thus computed as $f_{ND}=1-n_{\bar{H}}/n_{H}$, where $n_{\bar{H}}$ is the number of user-week pairs with a detected home, and $n_{H}$ is the number of user-week pairs with a ground truth home.
-In the case of D2, where ground truth information is not referring to a specific week, $n_{\bar{H}}$ corresponds to the number of users with a detected home, and $n_{H}$ corresponds to the number of users with a ground truth home.
-The “accuracy” captures the fraction of correctly labelled home locations among the set of detected home locations.
-Therefore, we define accuracy as $Acc=m_{\bar{H}}/n_{\bar{H}}$, where $m_{\bar{H}}$ is the number of user-weeks (for D1) or users (for D2) for which the detected home location matches the ground truth home location.
-In dataset D2, if annotators identified multiple homes for an individual, we considered it a match if the individual’s detected home location corresponded to at least one of the annotated ground truth home locations.
-Note that our definition does not correspond exactly to the traditional definition of accuracy because we separately assess the algorithm performance in terms of the detection/non-detection of a home (captured by $f_{ND}$) and the detection of the correct home (captured by $Acc$).
-The definitions above are also used in the case of work detection, where only users with a reported work location, $n_{W}$, are included in the denominator of $f_{ND}$, and only those with a detected and a matching work location are counted in the nominator and in the accuracy computation, as $n_{\bar{W}}$ and $m_{\bar{W}}$, respectively.
-5 Data and Code Availability
+8 Code and Data Availability
+All code needed to implement HoWDe is available as a GitHub repository at [61], along with usage examples in the GitHub folder /tutorials.
+We additionally provide HoWDe as a Python package on PyPI for direct installation via pip.
 Data to ensure replicability of the analysis performed in this work can be found at
-https://doi.org/10.11583/DTU.28846325 [Note: The link will become available upon publication].
+https://doi.org/10.11583/DTU.28846325.
 To protect user anonymity, we applied the following privacy-preserving transformations to the stop sequences.
-First, time was discretized into 10-minute intervals.
+First, time was discretised into 10-minute intervals.
 Second, each user’s timeline was shifted so that their first recorded activity occurred in the first week of 1970.
 Third, entire days were randomly shuffled within each week, while preserving the distinction between weekdays and weekends.
-All code needed to implement HoWDe is available at [37].
-We also provide HoWDe as a Python (PyPI) package at [54].
-Acknowledgments
-L.L.
-has been supported by the ERC project “IMMUNE” (Grant agreement ID: 101003183).
-L.L.
-acknowledges the support from the “Fondazione Romeo ed Enrica Invernizzi” for the research activities of the ‘Covid Crisis Lab’ at Bocconi University.
-L.A.
-and S.D.S.
-acknowledge support from the Independent Research Fund Denmark (DFF) through the DFF-Research Project 1 (Inge Lehmann) grant “Gender Gaps in Human Mobility”.
-The authors thank Y.
-Liao for providing thoughtful feedback on the code release, which helped improve its clarity and usability.
-Declaration of Interest statement
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
-Author Contributions
-S.D.S., L.L., S.F., and L.A.
-conceived the study.
-S.D.S., L.L., and O.D.L.C.
-curated and preprocessed the mobility and ground truth data.
-Two alternative algorithms were independently developed by L.L., S.F., and O.D.L.C., and by S.D.S.
-and L.A.
-The two designs were compared and discussed collaboratively, and the final HoWDe workflow is based on the design by S.D.S.
-and L.A., with L.L.
-contributing to its finalization.
-S.D.S.
-and L.L.
-designed the analysis pipeline and conducted the statistical analyses.
-S.F.
-and L.A.
-provided domain expertise and guidance on the interpretation of results.
-S.D.S.
-and L.L.
-drafted the manuscript.
-All authors discussed the results and approved the final version of the manuscript.
+All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).